### v3

# Borsuk-Ulam and Replicable Learning of Large-Margin Halfspaces

Abstract
Recent remarkable advances in learning theory have established that, for total concept classes, list replicability, global stability, differentially private (DP) learnability, and shared-randomness replicability all coincide with the finiteness of Littlestone dimension. Does this equivalence extend to partial concept classes?
We answer this question by proving that the list replicability number of $d$-dimensional $\gamma$-margin half-spaces satisfies
$\frac{d}{2}+1\leq\operatorname{LR}(\mathcal{H}^{d}_{\gamma})\leq d,$
which grows with dimension. Consequently, for partial classes, list replicability and global stability do not necessarily follow from bounded Littlestone dimension, pure DP-learnability, or shared-randomness replicability.
Applying our main theorem, we resolve several open problems:
  • Every disambiguation of infinite-dimensional large-margin half-spaces to a total concept class has unbounded Littlestone dimension, answering an open question of Alon, Hanneke, Holzman, and Moran (FOCS ’21).
  • The maximum list-replicability number of any finite set of points and homogeneous half-spaces in $d$-dimensional Euclidean space is $d$, resolving a problem of Chase, Moran, and Yehudayoff (FOCS ’23).
  • Every disambiguation of the Gap Hamming Distance problem in the large gap regime has unbounded public-coin randomized communication complexity. This answers an open question of Fang, Göös, Harms, and Hatami (STOC ’25).
  • There exists a partial concept class with Littlestone dimension $1$ such that all its disambiguations have infinite Littlestone dimension. This answers a question of Cheung, H. Hatami, P. Hatami, and Hosseini (ICALP ’23).
Our lower bound follows from a topological argument based on the local Borsuk-Ulam theorem of Chase, Chornomaz, Moran, and Yehudayoff (STOC ’24). For the upper bound, we construct a list-replicable learning rule using the generalization properties of SVMs.
1 Introduction
The large-margin half-space classification problem is a central topic in learning theory, extensively studied for its theoretical significance and practical impact. Theoretically, it is a fundamental and elegant geometric problem, and it serves as a simple model for understanding more complex learning problems. Practically, its importance stems from the success of Support Vector Machines (SVMs), which leverage the large-margin assumption to produce accurate classifications in high-dimensional spaces. SVMs play a key role in many real-world applications, including pattern and text recognition, image analysis, geographic information systems, bioinformatics, healthcare, and fraud detection.
To describe the large-margin classification problem, we first consider standard half-space classification without any margin assumption. After appropriate normalization and homogenization, we may assume that the domain is the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$, and the objective is to learn an unknown homogeneous half-space which classifies each point on the sphere as either ±$\pm 1$ depending on whether it belongs to the half-space.
The standard mathematical framework for analyzing the complexity of a learning task is probably approximately correct (PAC) learning. In this setting, the learner receives a training set of labeled points drawn independently from an unknown distribution $\mathcal{D}$, and the goal is to produce a hypothesis that accurately predicts the labels of new points sampled from $\mathcal{D}$ with high probability. The fundamental theorem of PAC learning states that the size of the training set required for PAC learning depends on a combinatorial parameter known as the VC dimension (Definition 2.5).
We also consider the Littlestone dimension (Definition 2.7), a refinement of the VC dimension that determines the optimal number of mistakes in online learning. In online learning, the learner receives data points sequentially and must predict the label of each point in turn before observing its correct label. The goal is to minimize the number of mistakes throughout the learning process.
Since the VC dimension of homogeneous half-spaces in $\mathbb{R}^{d}$ is $d$, PAC-learning this class requires a large training set in the high-dimensional settings. Moreover, the Littlestone dimension of this class is infinite except in the trivial case of $d=1$, meaning the class is not online learnable, even in $\mathbb{R}^{2}$.
Under the large-margin assumption, the learning task is restricted to points that lie at least a margin $\gamma>0$ away from the defining hyperplane. This constraint allows algorithms such as Perceptron [MP43, Ros58] and Support Vector Machines [VC64, CV95] to achieve efficient PAC and online learning, even in high-dimensional spaces. We study the large-margin halfspaces problem through the formal lens of partial classes, which offers a general framework for analyzing such constrained learning tasks.
A partial concept class over an arbitrary domain $\mathcal{X}$ is a set {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, where each $h\in\mathcal{H}$ is called a partial concept. The value $h(x)=\star$ indicates that $h$ is undefined at $x$, and therefore, both ±$\pm 1$ are acceptable predictions for the label of $x$.
In the large-margin setting, the domain is $\mathbb{S}^{d-1}$ and every homogeneous half-space defines a partial concept that assigns $h(x)=\star$ if $x$ lies within distance $\gamma$ of the defining hyperplane of $h$. Otherwise, it classifies $x$ as ±$\pm 1$ depending on whether it belongs to the half-space. More formally, each hypothesis $h_{w}:\mathbb{S}^{d-1}\to\{\pm 1,\star\}$ is specified by a unit vector $w\in\mathbb{S}^{d-1}$ and given by
$h_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}(\langle w,x\rangle)&\text{if }|\langle w,x\rangle|\geq\gamma\\ \star&\text{otherwise}\end{cases}.$ (1)
We denote the partial concept class of all such $h_{w}$ by $\mathcal{H}^{d}_{\gamma}$.
We also define $\mathcal{H}^{\infty}_{\gamma}$ as the class of all partial concepts $h_{w}:\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}\to\left\{\pm 1,\star\right\}$, where each hypothesis is specified by a unit vector $w\in\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}$ of arbitrary finite dimension. For any $x$ and $w$ of different dimensions, we set $h_{w}(x)\coloneqq\star$, and otherwise, define $h_{w}(x)$ as in Eq. 1.
In [AHHM22], Alon, Hanneke, Holzman, and Moran proved that, as in the case of total concept classes, the VC and Littlestone dimensions characterize PAC and online learnability of partial concept classes. In the case of $\mathcal{H}_{\gamma}^{d}$, the classic mistake-bound analysis of the Perceptron algorithm [MP43, Ros58] (see also [SSBD14, Theorem 9.1]) shows the following upper bound on the Littlestone and VC dimensions:
$\operatorname{VCdim}(\mathcal{H}_{\gamma}^{d})\leq\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})\leq\gamma^{-2}.$ (2)
Crucially, these bounds are independent of $d$, which explains the efficient PAC and online learnability of $\mathcal{H}_{\gamma}^{d}$ in arbitrarily high-dimensional spaces.
Recent advances in learning theory, sparked by the influential works on the differential privacy of PAC learning [ALMM19, BLM20, ABL+22], have established that for total concept classes, finite Littlestone dimension not only characterizes online learnability but is also equivalent to list replicability, global stability, differential privacy (DP) learnability, and shared-randomness replicability. We will discuss these notions in depth shortly. This naturally leads to the question of whether these equivalences extend to partial concept classes. The primary goal of this paper is to provide a negative answer to this question by proving that $\mathcal{H}_{\gamma}^{\infty}$ is not list replicable, even though it has a bounded Littlestone dimension by Eq. 2.
Before delving deeper into these results, we formally define some key notions. Consider a partial concept class {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$. A distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ is realizable by a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ if, for every $n$, a random sample $\bm{S}=((\bm{x}_{i},\bm{y}_{i}))_{i=1}^{n}\sim\mathcal{D}^{n}$ is almost surely realizable by some $h\in\mathcal{H}$, meaning that $h(\bm{x}_{i})=\bm{y}_{i}$ for all $i=1,\ldots,n$. Note that $\mathcal{D}$ is a distribution over $\mathcal{X}\times\left\{\pm 1\right\}$, so none of the labels $\bm{y}_{i}$ take the value $\star$.
Here, and throughout the paper, we use boldface letters to denote random variables and use the notation $(\bm{x},\bm{y})\sim\mathcal{D}$ to express that $(\bm{x},\bm{y})$ is a random variable distributed according to $\mathcal{D}$.
The population loss of a partial concept $h\in\left\{\pm 1,\star\right\}^{\mathcal{X}}$ with respect to a distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ is
$\mathcal{L}_{\mathcal{D}}(h)\coloneqq\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}[h(\bm{x})\neq\bm{y}].$
Throughout this work, a learning rule refers to a (randomized) function $\bm{\mathcal{A}}$ that maps any sample $S\in\bigcup_{n=0}^{\infty}(\mathcal{X}\times\left\{\pm 1\right\})^{n}$ to a hypothesis $\bm{\mathcal{A}}(S)\in\left\{\pm 1\right\}^{\mathcal{X}}$. Since our primary focus is sample complexity rather than computational efficiency, we impose no computability constraints on $\bm{\mathcal{A}}$.
Replicability.
Replicability refers to the ability of an algorithm or study to produce consistent results when repeated under similar conditions and with similar data. It is a key principle of the scientific method that ensures research findings are reliable and not due to chance.
In recent years, a rapidly growing body of research has emerged that introduced various rigorous formulations of replicability for learning algorithms, particularly in the context of the PAC learning framework [BLM20, MM22, CMY23, BGH+23, KVYZ23, EKK+23, EKM+23, MSS23, EHKS23, KKL+24, KKMV23]. We introduce only the notions that are central to our main results and refer the reader to prior works, such as [MSS23, KKMV23, MM22], for a discussion of other related notions. Throughout, we use replicability as an umbrella term referring to the various formulations of this general concept.
The concept of replicability in PAC learning first emerged in [BLM20, ABL+22] in the study of differential privacy of PAC learning algorithms. These works introduced a notion of replicability known as global stability to derive privacy guarantees from online learnability.
A learning rule $\bm{\mathcal{A}}$ for a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is $(\epsilon,\rho)$-globally stable if for every realizable distribution $\mathcal{D}$, there is a hypothesis $h\in\left\{\pm 1\right\}^{\mathcal{X}}$ with population loss $\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon$ satisfying
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})=h]\geq\rho,\ \ \text{where }n=n(\epsilon).$
In other words, when run with samples from $\mathcal{D}$, there is a non-negligible chance $\rho>0$ that the learner will output the same hypothesis $h$. We define $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})$ to be the supremum of $\rho$ such that there is a $(\epsilon,\rho)$-globally stable learner for $\mathcal{H}$. The global stability parameter of $\mathcal{H}$ is then defined as
$\rho^{\mathrm{gs}}(\mathcal{H})\coloneqq\inf_{\epsilon>0}\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H}).$
The definition of global stability might initially seem weak, as a globally stable learner is not necessarily a PAC learner. In particular, since $\rho$ can be a small constant, there may be a probability as great as $1-\rho$ that the learning rule outputs a hypothesis with large population loss. However, [CMY23] showed that for total classes {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$, global stability is equivalent to a stronger notion called list replicability.
Definition 1.1 (List replicability).
A learning rule $\bm{\mathcal{A}}$ with sample complexity $n(\epsilon,\delta)$ is said to be an $(\epsilon,L)$-list replicable learner for {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$ if the following holds. For every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$, there is a list of total hypotheses $h_{1},\ldots,h_{L}\in\left\{\pm 1\right\}^{\mathcal{X}}$ such that
  • The population loss of every $h_{i}$ with respect to $\mathcal{D}$ is at most $\epsilon$:
    $\mathcal{L}_{\mathcal{D}}(h_{i})\leq\epsilon\text{ for every }i=1,\ldots,L.$
  • The probability that $\bm{\mathcal{A}}$ outputs a hypothesis outside the list is at most $\delta$:
    $\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})\not\in\left\{h_{1},\ldots,h_{L}\right\}]\leq\delta\ \text{ where }n=n(\epsilon,\delta).$
Given an accuracy parameter $\epsilon\in(0,1)$, the $\epsilon$-list replicability number of $\mathcal{H}$, denoted by $\operatorname{LR}_{\epsilon}(\mathcal{H})$, is the smallest integer $L$, such that there exists an $(\epsilon,L)$-list replicable learner for $\mathcal{H}$. We define $\operatorname{LR}_{\epsilon}(\mathcal{H})=\infty$ if no such integer exists.
The list replicability number of $\mathcal{H}$ is $\operatorname{LR}(\mathcal{H})\coloneqq\sup_{\epsilon>0}\operatorname{LR}_{\epsilon}(\mathcal{H}).$ We say $\mathcal{H}$ is list replicable if $\operatorname{LR}(\mathcal{H})<\infty$.
Definition 1.1 provides a strong notion of replicability as the learner’s output is typically chosen from a small list $\left\{h_{1},\ldots,h_{L}\right\}$, and all these hypotheses have small population loss.
In [CMY23], Chase, Moran and Yehudayoff proved that for every total class {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$, list replicability is equivalent to global stability. It is easy to check that their proof applies to partial concept classes, resulting in the following relationship between the list replicability number and global stability parameter.
Theorem 1.2 ([CMY23]).
For every partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, we have
$\rho^{\mathrm{gs}}(\mathcal{H})=\frac{1}{\operatorname{LR}(\mathcal{H})}.$
For completeness, we provide a proof in Appendix A.
As shown in [BLM20], the global stability parameter of any nontrivial concept class is at most $1/2$. This fundamental limitation implies that no nontrivial class can achieve true replicability in the strongest sense, where the learning algorithm outputs the same hypothesis with high probability. To remedy this, Impagliazzo, Lei, Pitassi, and Sorrell [ILPS22] introduced a weaker notion of replicability where different executions of the algorithm can use the same random seed. Formally, let $\mathcal{A}(S,r)$ be a randomized learning rule, where $r$ denotes the random seed.
Definition 1.3 (Shared-randomness replicability).
A concept class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$ is shared-randomness replicable if there exists a learning rule $\mathcal{A}$ and a sample complexity function $n(\epsilon,\delta)$ such that, for every $\epsilon,\delta>0$ and every realizable distribution $\mathcal{D}$, the following conditions hold:
  • Small population loss: $\Pr_{\bm{S}\sim\mathcal{D}^{n},\bm{r}}[\mathcal{L}_{\mathcal{D}}\left(\mathcal{A}(\bm{S},\bm{r})\right)>\epsilon]\leq\delta.$
  • Replicability with shared randomness: $\Pr_{\bm{S,S^{\prime}}\sim\mathcal{D}^{n},\bm{r}}[\mathcal{A}(\bm{S},\bm{r})=\mathcal{A}(\bm{S^{\prime}},\bm{r})]\geq 1-\delta.$
The following theorem, which brings together several celebrated recent results in learning theory, establishes that for total concept classes, all these notions of replicability, along with approximate differential privacy, are characterized by the finiteness of the Littlestone dimension.
Theorem 1.4 ([ALMM19, BLM20, ABL+22, CMY23, ILPS22]).
Let {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$ be a total concept class. The following statements are equivalent.
  • $\operatorname{Ldim}(\mathcal{H})<\infty$.
  • $\mathcal{H}$ is globally stable and equivalently list replicable.
  • $\mathcal{H}$ is shared-randomness replicable.
  • $\mathcal{H}$ is approximately differentially private (DP)-learnable (See Definition 2.9).
1.1 Main contributions
Replicability of large-margin half-spaces.
The differentially private and replicable learning of large-margin half-spaces have been studied extensively [BDMN05, LNUZ20, BMNS19, KMST20, BCS20, BMS22a, BMS22b, ILPS22, KKL+24].
In [LNUZ20], Nguyen, Ullman, and Zakynthinou gave a DP-learner for $\mathcal{H}^{d}_{\gamma}$ with a sample complexity that does not depend on the dimension $d$. In fact, they proved a stronger result that $\mathcal{H}^{d}_{\gamma}$ is pure DP-learnable (Definition 2.10). Combining the dimension-free DP-learnability of $\mathcal{H}^{d}_{\gamma}$ with the “DP-learnable to Shared-randomness Replicability” reduction of [BGH+23] shows that $\mathcal{H}^{d}_{\gamma}$ is shared-randomness replicable with dimension-free sample complexity. More recently, Kalavasis, Karbasi, Larsen, Velegkas, and Zhou [KKL+24] employed a more sophisticated rounding scheme to directly obtain a shared-randomness replicability of this problem, which further improves the sample complexity. In the following theorem, we summarize the consequences of these lines of work that are relevant to this paper.
Theorem 1.5 ([MP43, Ros58, LNUZ20, KKL+24]).
For every $\gamma>0$ and $d\in\mathbb{N}\cup\left\{\infty\right\}$, the class $\mathcal{H}^{d}_{\gamma}$ satisfies the following.
  • $\operatorname{Ldim}(\mathcal{H}^{d}_{\gamma})<\gamma^{-2}$.
  • The class $\mathcal{H}^{d}_{\gamma}$ is (pure) DP-learnable with dimension-independent sample complexity.
  • The class $\mathcal{H}^{d}_{\gamma}$ is shared-randomness replicable with dimension-independent sample complexity.
Given the positive results in Theorem 1.5 and the equivalences in Theorem 1.4, one might expect that, as in the case of total concepts, the list replicability number of $\mathcal{H}^{d}_{\gamma}$ is independent of $d$. Our main theorem establishes that this is not the case—the list replicability number of $\mathcal{H}^{d}_{\gamma}$ grows as $d$ increases.
Theorem 1.6 (Main Theorem).
For any fixed margin $\gamma\in(0,1)$, finite dimension $d>1$, and accuracy parameter $\epsilon\in(0,1/2)$,
$\frac{d}{2}+1\leq\operatorname{LR}_{\epsilon}(\mathcal{H}_{\gamma}^{d})\leq d.$
In particular, $\mathcal{H}_{\gamma}^{\infty}$ is not list replicable (equiv. globally stable).
Theorem 1.6 reveals a surprising distinction in the partial setting: list replicability does not follow from bounded Littlestone dimension, replicability, DP-learnability, or even pure DP-learnability. Specifically, for every $\gamma\in(0,1)$, the partial concept class $\mathcal{H}^{\infty}_{\gamma}$ is shared-randomness replicable, pure DP-learnable, and has a finite Littlestone dimension, but it is not list replicable.
The lower bound in Theorem 1.6 relies on a topological argument involving covers of the sphere by antipodal-free open sets. In particular, we apply the local Borsuk-Ulam theorem of [CCMY24], which states that in such a cover, there is a point that belongs to at least $\frac{d}{2}+1$ sets. Alternatively, one could use Ky Fan’s classical theorem [Fan52], but this would yield the slightly weaker bound of $\frac{d}{2}$.
For the upper bound, we construct a learning rule that uses the generalization properties of hard-SVM combined with a list-replicable rounding scheme using a fine net in general position.
Disambiguations of Large-margin Half-spaces:
One of the central questions studied in [AHHM22] is whether the learnability of a partial concept class can always be inferred from the learnability of a suitable extension of it to a total concept class.
Formally, a disambiguation of a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is a total concept class ¯{±$\overline{\mathcal{H}}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$ such that for every $h\in\mathcal{H}$ and every finite $S\subseteq h^{-1}(\left\{\pm 1\right\})$, there exists an $\bar{h}\in\overline{\mathcal{H}}$ that is consistent with $h$ on $S$. Intuitively, one may think of disambiguation as replacing each $\star$ with $-1$ or $+1$, although this intuition is not completely rigorous in the infinite case.
Note that any distribution realizable by a partial concept class $\mathcal{H}$ is also realizable by any disambiguation $\overline{\mathcal{H}}$ of $\mathcal{H}$. Consequently, if a disambiguation $\overline{\mathcal{H}}$ is efficiently learnable, the same holds true for $\mathcal{H}$.
Motivated by the fact that $\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})\leq\gamma^{-2}$, [AHHM22] posed the following question.
Question 1.7 ([AHHM22, Question 4]).
Does there exist a disambiguation of $\mathcal{H}_{\gamma}^{d}$ by a total concept class whose VC/Littlestone dimension is bounded by a function of $\gamma$?
In the following theorem, we provide a negative answer to the Littlestone dimension part of 1.7.
Theorem 1.8.
For every $d\in\mathbb{N}$, every disambiguation $\overline{\mathcal{H}}$ of $\mathcal{H}_{\gamma}^{d}$ satisfies $\operatorname{Ldim}(\overline{\mathcal{H}})=\Omega(\sqrt{\log d})$.
Proof.
This follows by combining the lower bound of Theorem 1.6 and a result of [GGKM21] showing that for every total concept class $\mathcal{H}$, $\operatorname{LR}_{\epsilon}(\mathcal{H})\leq 2^{O_{\epsilon}\left(\operatorname{Ldim}(\mathcal{H})^{2}\right)}$. ∎
On the other hand, there is a disambiguation $\overline{\mathcal{H}}$ of $\mathcal{H}_{\gamma}^{d}$ with Littlestone dimension $\operatorname{Ldim}(\overline{\mathcal{H}})=O(d)$ since [HHM23] shows that there is a disambiguation of $\mathcal{H}_{\gamma}^{d}$ into a finite total concept class $\overline{\mathcal{H}}$ with $2^{O(d)}$ concepts. The upper bound on the Littlestone dimension follows from the fact that every finite class $\mathcal{H}$ satisfies $\operatorname{VCdim}(\mathcal{H})\leq\operatorname{Ldim}(\mathcal{H})\leq\log|\mathcal{H}|$.
Theorem 1.8 has two important implications. Firstly, it shows $\mathcal{H}^{\infty}_{\gamma}$ is a partial class with finite Littlestone dimension, but any disambiguation of it has an infinite Littlestone dimension. The existence of such partial classes was posed as a question in [AHHM22] and answered in [CHHH23] through a complex construction based on Göös’ breakthrough [Göö15] in communication complexity. Theorem 1.8 offers a more natural example of a concept class with this property.
Secondly, the construction of  [CHHH23] has Littlestone dimension $2$, and in [CHHH23, Question 4.1], they asked whether there exists a partial concept class with Littlestone dimension $1$ such that all its disambiguations have infinite Littlestone dimension. The following corollary of Theorem 1.8 resolves their question.
Corollary 1.9.
Let $\gamma\in(1/\sqrt{2},1)$. Then $\operatorname{Ldim}(\mathcal{H}_{\gamma}^{\infty})=1$, while any disambiguation of $\mathcal{H}_{\gamma}^{\infty}$ has infinite Littlestone dimension.
Proof.
We have $21\leq\operatorname{Ldim}(\mathcal{H}_{\gamma}^{\infty})\leq 1/\gamma^{2}<$ 2. An application of Theorem 1.8 completes the proof. ∎
List replicability of finite classes.
In [CMY23, Theorems 5 and 13], Chase, Moran, and Yehudayoff proved that the list replicability number of any finite set of homogeneous half-spaces and points that do not lie on the half-spaces in $\mathbb{R}^{2}$ is at most $2$. They further asked whether a similar finite bound holds in higher dimensions. As we discuss below, we give an affirmative answer to their question in Theorem 1.11.
Definition 1.10.
Define the finitary list replicability number of a concept class {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$, denoted $\widetilde{\operatorname{LR}}(\mathcal{H})$, as
$\widetilde{\operatorname{LR}}(\mathcal{H})\coloneqq\sup_{\text{finite }S\subseteq\mathcal{X}}\operatorname{LR}(\mathcal{H}|_{S}).$
Given $d\in\mathbb{N}$, let $\mathcal{H}^{d}$ be the partial concept class defined by points in $\mathbb{S}^{d-1}$ and partial concepts $h_{w}:\mathbb{S}^{d-1}\to\left\{\pm 1,\star\right\}$, where $w\in\mathbb{S}^{d-1}$, given by
$h_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}(\langle w,x\rangle)&\text{if }\langle w,x\rangle\neq 0\\ \star&\text{otherwise}\end{cases}.$
Note that for every $\gamma>0$, $\mathcal{H}^{d}_{\gamma}$ is a subclass of $\mathcal{H}^{d}$.
Since $\mathcal{H}^{2}$ has infinite threshold dimension [AHHM22, Definition 28], by the celebrated result of [ABL+22], it is not globally stable and therefore $\operatorname{LR}(\mathcal{H}^{2})=\infty$. Consequently, the aforementioned result of [CMY23] implies the following interesting gap:
$\operatorname{LR}(\mathcal{H}^{2})=\infty\qquad\text{while}\qquad\widetilde{\operatorname{LR}}(\mathcal{H}^{2})=2.$
The authors of [CMY23] ask whether a similar bound holds for $\widetilde{\operatorname{LR}}(\mathcal{H}^{3})$. In the following theorem, we answer their question by extending their result to all dimensions.
Theorem 1.11.
For every dimension $d>1$, we have $\widetilde{\operatorname{LR}}(\mathcal{H}^{d})=d$.
Proof.
The upper bound is an easy consequence of the upper bound of our main theorem (Theorem 1.6). Indeed, any finite set of points $S\subset\mathbb{S}^{d-1}$ and hypotheses $H\subset\mathcal{H}^{d}|_{S}$ defined by unit vectors $W\subset\mathbb{S}^{d-1}$ is a sub-concept class of $\mathcal{H}^{d}_{\gamma}$, where $\gamma=\min_{x\in S,w\in W}|\langle w,x\rangle|$. The claim now follows from our upper bound from Theorem 1.6.
For the lower bound, we use a result of Chase, Moran, and Yehudayoff [CMY23, Theorem 3] stating that for every concept class $\mathcal{H}$,
$\operatorname{LR}(\mathcal{H})\geq\operatorname{VCdim}(\mathcal{H}).$
The result now follows as $\operatorname{VCdim}(\mathcal{H}^{d})=d$, and therefore, $\mathcal{H}^{d}$ has a finite subclass of VC dimension $d$. ∎
The sign-rank of a sign matrix $A$, denoted by ${\mathbf{rk}}_{\pm}(A)$, is the smallest rank of a real matrix $B_{\mathcal{X}\times\mathcal{Y}}$ such that the entries of $B$ are nonzero and have the same signs as their corresponding entries in $A$. The definition of sign-rank naturally extends to partial matrices, where for invalid entries of $A$, the corresponding entry in $B$ could be any real number.
Geometrically, sign-rank is the smallest dimension in which the matrix is realized as points and homogeneous half-spaces. We will state the definition of sign-rank in this terminology for partial classes.
Definition 1.12 (Sign-rank).
The sign-rank of a partial class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, denoted by ${\mathbf{rk}}_{\pm}(\mathcal{H})$, is the smallest $d$ such that there exist vectors $u_{h},v_{x}\in\mathbb{R}^{d}$ for all pairs $h\in\mathcal{H},x\in\mathcal{X}$ such that $h(x)=\operatorname{sgn}({\left\langle u_{h},v_{x}\right\rangle})$ whenever $h(x)\neq\star$.
Theorem 1.11 and the VC lower bound of [CMY23, Theorem 3] immediately imply the following general bounds on $\widetilde{\operatorname{LR}}(\mathcal{H})$.
Corollary 1.13.
For every partial class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, we have
$\operatorname{VCdim}(\mathcal{H})\leq\widetilde{\operatorname{LR}}(\mathcal{H})\leq\operatorname{rank_{\pm}}(\mathcal{H}).$
Gap Hamming Distance.
The discrete analogue of large-margin half-spaces is the well-studied Gap Hamming Distance (GHD) problem, a central problem in communication complexity. For $n\in\mathbb{N}$ and $\gamma\in(0,1)$, the $n$-bit $\operatorname{GHD}_{\gamma}$ problem, denoted $\operatorname{GHD}_{\gamma}^{n}$, is the partial function on inputs $x,y\in\{\pm 1\}^{n}$ defined by
$\operatorname{GHD}_{\gamma}^{n}(x,y)\coloneqq\begin{cases}\operatorname{sgn}({\left\langle x,y\right\rangle})&\text{if $|{\left\langle x,y\right\rangle}|\geq\gamma n$}\\ \star&\text{otherwise}\end{cases}.$
For fixed $\gamma\in(0,1)$, the public-coin randomized communication complexity of $\operatorname{GHD}_{\gamma}$ is at most a fixed constant depending only on $\gamma$. This places $\operatorname{GHD}_{\gamma}$ in the complexity class $\sf{BPP}_{0}$, which consists of communication problems with constant-cost public-coin randomized protocols. Moreover, $\operatorname{GHD}_{\gamma}$ exhibits certain properties that separate $\sf{BPP}_{0}$ from some complexity classes (e.g., $\sf{UPP}_{0}$) and measures [HHH23, CLV19, HHM23, FGHH25, Son14].
A key caveat of these separations is that $\operatorname{GHD}_{\gamma}$ is a partial function, and whether it extends to any family of total functions with similar properties remained an open problem before this work. One suggested approach [FGHH25] to resolving these questions is to find a disambiguation of $\operatorname{GHD}_{\gamma}$ to a total function with constant public-coin randomized communication complexity. In Corollary 1.15 below, we prove this is impossible: every disambiguation of $\operatorname{GHD}_{\gamma}^{n}$ has public-coin randomized communication complexity at least $\Omega(\log\log n)$.
First, we show that every such disambiguation has an unbounded Littlestone dimension.
Theorem 1.14.
Let $\gamma\in(0,1)$ be a margin parameter. Every family of disambiguations $\{M_{n}\}_{n=1}^{\infty}$ of the Gap Hamming Distance matrices $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$ satisfies
$\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n}).$
We obtain the following corollary by combining Theorem 1.14 with some known results.
Corollary 1.15.
Let $\gamma\in(0,1)$ be a margin parameter. Every family of disambiguation of the Gap Hamming Distance matrices $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$ has public-coin randomized communication complexity $\Omega(\log\log n)$.
The bound follows by combining Theorem 1.14 with the known relationship between Littlestone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity. See Section 4 for the proof.
1.2 Concluding remarks and open problems
While Theorem 1.4 provides an elegant and comprehensive characterization of various notions of replicability and privacy for total concepts through the combinatorial framework of the Littlestone dimension, the landscape of replicability for partial classes, as demonstrated by the results of this paper, is more intricate and less understood.
The “DP-learnability to Shared-randomness Replicability” reduction from [BGH+23] extends to the partial setting. Moreover, [FHM+24] recently showed that for partial classes, DP-learnability implies a finite Littlestone dimension.
Theorem 1.16 ([BGH+23, FHM+24]).
Let {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class.
  • If $\mathcal{H}$ is DP-learnable, then $\mathcal{H}$ is shared-randomness replicable.
  • If $\mathcal{H}$ is DP-learnable, then $\operatorname{Ldim}(\mathcal{H})<\infty$.
On the other hand, our main theorem shows that for partial concepts, list replicability does not follow from bounded Littlestone dimension, shared-randomness replicability, DP-learnability, or even pure DP-learnability. It is also straightforward to show that, even in the partial setting, list replicability implies shared-randomness replicability.
Proposition 1.17 ([KKMV23, Lemma 8]).
If {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is list replicable (equiv. globally stable), then $\mathcal{H}$ is shared-randomness replicable.
To our knowledge, no further relationships among the four notions of DP-learnability, shared-randomness replicability, Littlestone dimension, and list replicability are known for partial functions.
Open problems.
Finally, we list some open problems for future research that naturally arise from our work.
  1. 1.
    Are DP-learnability, shared-randomness replicability, and finite Littlestone dimension equivalent for partial functions? If not, what are the precise relationships between them?
  2. 2.
    Is there a combinatorial notion of dimension that characterizes list replicability?
  3. 3.
    How tight are the inequalities in Corollary 1.13, namely,
    $\operatorname{VCdim}(\mathcal{H})\leq\widetilde{\operatorname{LR}}(\mathcal{H})\leq\operatorname{rank_{\pm}}(\mathcal{H})?$
  4. 4.
    Is the answer to 1.7 negative for VC dimension?
    In an article posted on arXiv just a few days ago, Chornomaz, Moran, and Waknine [CMW25] explored this problem using a topological approach, but the question remains open.
2 Preliminaries
Much of the notation and definitions were already outlined in the introduction. In this section, we review a few key concepts from learning theory, specifically PAC learnability, the VC and Littlestone dimensions, and differential privacy.
In some of our proofs, we will apply Hoeffding’s classic concentration inequality, which we state below for reference.
Theorem 2.1 (Hoeffding’s inequality).
Let $c\in\mathbb{R}$ and let $\bm{x}_{1},\ldots,\bm{x}_{n}$ be independent random variables with $\bm{x}_{i}\in[-c,c]$ and $\mathbb{E}[\bm{x}_{i}]=0$. Then, for any $t>0$,
$\Pr\left[\left|\sum_{i=1}^{n}\bm{x}_{i}\right|\geq t\right]\leq 2e^{-\frac{t^{2}}{2c}}.$
We define the population loss of a partial concept $h\in\left\{\pm 1,\star\right\}^{\mathcal{X}}$ with respect to a distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ as
$\mathcal{L}_{\mathcal{D}}(h)\coloneqq\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}[h(\bm{x})\neq\bm{y}].$
Under this definition, predictions with $h(x)=\star$ always count as mistakes. The population loss of a class $\mathcal{H}$ is given by the smallest population loss achievable within $\mathcal{H}$:
$\mathcal{L}_{\mathcal{D}}(\mathcal{H})\coloneqq\inf_{h\in\mathcal{H}}\mathcal{L}_{\mathcal{D}}(h).$
The following simple lemma from [AHHM22] establishes the connection between realizability and having zero population loss.
Lemma 2.2 ([AHHM22]).
Let {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class, and let $\mathcal{D}$ be a distribution on $\mathcal{X}\times\left\{\pm 1\right\}$. If $\mathcal{L}_{\mathcal{D}}(\mathcal{H})=0$, then $\mathcal{D}$ is realizable by $\mathcal{H}$. Conversely, if $\mathcal{D}$ is realizable and has finite or countable support, then $\mathcal{L}_{\mathcal{D}}(\mathcal{H})=0$.
PAC learning of partial classes:
Given a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, the goal of PAC (Probably Approximately Correct) learning is for the learning rule to produce, with high probability, a hypothesis whose population loss is close to the best achievable loss within $\mathcal{H}$.
Definition 2.3 (PAC learning of partial concept classes).
A class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is PAC learnable if there is a learning rule $\bm{\mathcal{A}}$ and a function $n(\epsilon,\delta)$ such that for every $\epsilon,\delta>0$ and every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm\right\}$, we have
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\mathcal{L}_{\mathcal{D}}(\bm{\mathcal{A}}(\bm{S}))\leq\epsilon]\geq 1-\delta\ \text{ where }n=n(\epsilon,\delta).$
2.1 VC and Littlestone dimensions
We now extend the definition of the VC dimension to partial concept classes.
Definition 2.4 (Shattered set).
A finite set of points $C=\left\{x_{1},\ldots,x_{n}\right\}\subseteq\mathcal{X}$ is shattered by a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ if, for every pattern $y\in\left\{\pm 1\right\}^{n}$, there exists $h\in\mathcal{H}$ such that $h(x_{i})=y_{i}$ for all $i\in[n]$.
Definition 2.5 (VC dimension).
The VC dimension of a partial concept class $\mathcal{H}$, denoted $\operatorname{VCdim}(\mathcal{H})$, is the largest integer $d$ for which there exists a size-$d$ subset of $\mathcal{X}$ that is shattered by $\mathcal{H}$. If no such maximum $d$ exists, we define $\operatorname{VCdim}(\mathcal{H})=\infty$.
Viewed as a matrix, the VC dimension of $\mathcal{H}$ is the largest $d$ such that the associated partial matrix $M_{\mathcal{X}\times\mathcal{H}}$ contains a $d\times 2^{d}$ submatrix with ±$\pm 1$ entries, where the columns enumerate all $d$-bit ±$\pm 1$ patterns.
The fundamental theorem of PAC learning asserts that a total concept class is PAC learnable if and only if its VC dimension is finite. This result was extended to partial concept classes by [AHHM22] using a completely different proof.
Theorem 2.6 ([AHHM22]).
The following statements are equivalent for any partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$.
  • $\operatorname{VCdim}(\mathcal{H})<\infty$;
  • $\mathcal{H}$ is PAC learnable;
Littlestone dimension:
In his influential work, Littlestone [Lit88] provided a combinatorial characterization of concept classes that are online learnable. To this end, he introduced a measure that extends the VC dimension by considering the shattering of decision trees rather than sets.
A mistake tree of depth $d$ over a domain $\mathcal{X}$ is a complete binary tree of depth $d$ with the following properties:
  • Each internal node is labeled by an element $x\in\mathcal{X}$.
  • Each edge is labeled by a binary value $b\in\left\{\pm 1\right\}$, where $b=-1$ indicates a left child and $b=1$ indicates a right child.
Every root-to-leaf path in the tree is described by a sequence
$(x_{1},b_{1}),\ldots,(x_{d},b_{d}),$
where $x_{i}\in\mathcal{X}$ is the label of the $i$th internal node along the path, and $b_{i}$ specifies whether the path moves to the left or right child at each level.
A concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ shatters a mistake tree if, for every root-to-leaf path
$(x_{1},b_{1}),\ldots,(x_{d},b_{d}),$
there exists a hypothesis $h\in\mathcal{H}$ such that $h(x_{i})=b_{i}$ for all $i\in[d]$.
Definition 2.7 (Littlestone Dimension).
The Littlestone dimension of a concept class $\mathcal{H}$, denoted $\operatorname{Ldim}(\mathcal{H})$, is the largest integer $d$ for which there exists a mistake tree of depth $d$ that is shattered by $\mathcal{H}$. If no such maximum $d$ exists, we define $\operatorname{Ldim}(\mathcal{H})=\infty$.
It always holds that $\operatorname{VCdim}(\mathcal{H})\leq\operatorname{Ldim}(\mathcal{H})$, since any set $C=\left\{x_{1},\ldots,x_{d}\right\}$ shattered by $\mathcal{H}$ gives rise to a mistake tree of depth $d$, where all nodes at level $i$ are labeled with $x_{i}$. This tree is shattered by $\mathcal{H}$.
Littlestone proved that a total concept class $\mathcal{H}$ is online learnable if and only if $\operatorname{Ldim}(\mathcal{H})<\infty$. This result was later extended to partial concept classes by [AHHM22].
2.2 Differentially Private Learnability
The widely adopted approach for ensuring privacy in machine learning is the differential privacy (DP) framework, introduced in [DMNS06]. Informally, differential privacy in learning means that no single labeled example in the input dataset significantly impacts the learner’s output hypothesis. In other words, the output distribution of a differentially private randomized learning algorithm remains nearly unchanged if a single data point is modified.
Differential privacy is quantified with two parameters $\epsilon,\delta>0$. We say that two probability distributions $p$ and $q$ are $(\epsilon,\delta)$-indistinguishable, if for every event $E$, we have
$p(E)\leq e^{\epsilon}q(E)+\delta\ \text{ and }\ q(E)\leq e^{\epsilon}p(E)+\delta.$
Two random variables are $(\epsilon,\delta)$-indistinguishable if their distributions satisfy this condition.
Definition 2.8 (Differential Privacy).
Given $\epsilon,\delta>0$, a randomized learning rule
$\bm{\mathcal{A}}:(\mathcal{X}\times\left\{\pm 1\right\})^{n}\to\left\{\pm\right\}^{\mathcal{X}}$
is $(\epsilon,\delta)$-differentially-private if for every two samples $S,S^{\prime}\in(\mathcal{X}\times\left\{\pm\right\})^{n}$ differing on a single example, the random variables $\bm{\mathcal{A}}(S)$ and $\bm{\mathcal{A}}(S^{\prime})$ are $(\epsilon,\delta)$-indistinguishable.
We emphasize that $(\epsilon,\delta)$-indistinguishability must hold for every such pair of samples, regardless of whether they are drawn from a (realizable) distribution.
The special case where $\delta=0$ is known as pure differential privacy, while the more general case where $\delta>0$ is referred to as approximate differential privacy.
In approximate differential privacy, the parameters $\epsilon$ and $\delta$ are typically set as follows: $\epsilon$ is taken to be a small fixed constant (e.g., 0.1), while $\delta$ is a negligible function, $\delta=n^{\omega(1)}$, where $n$ is the sample size. Definitions 2.9 and 2.10 formally define differentially private learnability under the assumption that the privacy parameter $\epsilon$ is fixed at $\epsilon=0.1$.
Definition 2.9 (Approximate Differentially Private Learnability).
We say that a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is approximate differentially private learnable (DP-learnable) if there is a learning rule $\bm{\mathcal{A}}:(\mathcal{X}\times\left\{\pm 1\right\})^{*}\to\left\{\pm 1\right\}^{\mathcal{X}}$ with sample complexity $n(\epsilon,\delta,\delta^{\prime})$ such that the following holds.
  • For every $\epsilon,\delta,\delta^{\prime}>0$, the class $\mathcal{H}$ is $(\epsilon,\delta)$-PAC learnable by $\bm{\mathcal{A}}$ using $n(\epsilon,\delta,\delta^{\prime})$ samples.
  • For every $\epsilon,\delta,\delta^{\prime}>0$, the learning rule $\bm{\mathcal{A}}$ applied to samples of size $n(\epsilon,\delta,\delta^{\prime})$ is $(0.1,\delta^{\prime})$-differentially private.
  • For every fixed $\epsilon,\delta>0$, we have
    $\lim_{\delta^{\prime}\to 0}\frac{\log(1/\delta^{\prime})}{\log n(\epsilon,\delta,\delta^{\prime})}=\infty.$
Definition 2.10 (Pure Differentially Private Learnability).
We say that a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is pure differentially private learnable (pure DP-learnable) if $\mathcal{H}$ is PAC learnable by a $(0.1,0)$-differentially private learning rule.
3 Proof of the Main Theorem
3.1 The Lower Bound
We prove the lower bound via a topological argument that utilizes the following local version of the Borsuk-Ulam theorem proved in [CCMY24].
Theorem 3.1 (Local Borsuk-Ulam [CCMY24]).
Let $d\geq 2$ be an integer. If $\mathcal{F}$ is a finite antipodal-free open cover of the sphere $\mathbb{S}^{d-1}$, then there exists some $w\in\mathbb{S}^{d-1}$ contained in at least $\lceil\frac{d}{2}+1\rceil$ member sets of $\mathcal{F}$.
Fix any margin $\gamma\in(0,1)$, dimension $d\geq 2$ and $\epsilon\in(0,1/2)$, and suppose that $\bm{\mathcal{A}}$ is an $(\epsilon,L)$-list replicable learning rule for $\mathcal{H}_{\gamma}^{d}$. We prove that $L\geq\frac{d}{2}+1$.
By the definition of list replicability, for any $\delta>0$, there is an integer $n$ so that for any realizable distribution $\mathcal{D}$, there exists a list of hypotheses $\{h_{1},\dots,h_{L}\}$ with
$\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\mathbf{S})\in\{h_{1},\dots,h_{L}\}]\geq 1-\delta\text{ and }\mathcal{L}_{\mathcal{D}}(h_{i})\leq\epsilon\text{ for all }i\in[L].$
Now pick any $\alpha>0$ and $\epsilon^{\prime}\in(\epsilon,1/2)$. By taking $\delta$ sufficiently small, for any distribution $\mathcal{D}$, we can choose a hypothesis $h_{\mathcal{D}}\in\{h_{1},\dots,h_{L}\}$ such that
$\Pr_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{\mathcal{D}}]>\frac{1}{L+\alpha}\text{ and }\mathcal{L}_{\mathcal{D}}(h_{\mathcal{D}})<\epsilon^{\prime}.$ (3)
We will focus on a certain collection of realizable distributions $\mathcal{D}$ on $\mathbb{S}^{d-1}\times\{\pm 1\}$. For any $w\in\mathbb{S}^{d-1}$, take $\mathcal{D}_{w}$ to be the uniform distribution on the set $\{(x,h_{w}(x))\ |\ x\in{\mathrm{supp}}(h_{w})\}$. These distributions are, by definition, realizable. Hence, for each $\mathcal{D}_{w}$, we can choose some particular hypothesis $h_{\mathcal{D}_{w}}$ that satisfies the conditions in Eq. 3. Collect these hypotheses in a set $T$, that is to say
$T\coloneqq\{h_{\mathcal{D}_{w}}\ |\ w\in\mathbb{S}^{d-1}\}.$
For each $h\in T$, define the set $C_{h}\subset\mathbb{S}^{d-1}$ as
$C_{h}\coloneqq\left\{w\in\mathbb{S}^{d-1}\ |\ \Pr_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h]>\frac{1}{L+\alpha}\text{ and }\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}\right\}.$
Claim 3.2.
The family $\{C_{h}\}_{h\in T}$ forms an antipodal-free open cover of $\mathbb{S}^{d-1}$.
Proof.
The fact that any set $C_{h}$ is antipodal-free follows from the accuracy constraint $\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}$. Indeed, for any $w\in\mathbb{S}^{d-1}$, the concepts $h_{w}$ and $h_{-w}$ have identical support, on which they disagree at every point. Thus the population loss of any hypothesis $h$ satisfies the equation
$\mathcal{L}_{\mathcal{D}_{w}}(h)+\mathcal{L}_{\mathcal{D}_{-w}}(h)=1.$
For any $w\in C_{h}$, we have that $\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}<1/2$, whereby $w$ and $-w$ cannot both be in $C_{h}$.
Next, each set $C_{h}$ is open because both $\Pr_{\bm{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\bm{S})=h]$ and $\mathcal{L}_{\mathcal{D}_{w}}(h)$ are continuous in $w$. Lastly, the family $\{C_{h}\}_{h\in T}$ covers $\mathbb{S}^{d-1}$ because, for any $w\in\mathbb{S}^{d-1}$, the set $C_{h_{\mathcal{D}_{w}}}$ contains $w$ by construction. ∎
Now note that the antipodal-free open cover $\{C_{h}\}_{h\in T}$ admits a finite subcover by the compactness of the unit sphere. Applying Theorem 3.1 to such a finite subcover guarantees that some $w\in\mathbb{S}^{d-1}$ is contained in at least $t\coloneqq\lceil\frac{d}{2}+1\rceil$ sets $C_{h_{1}},C_{h_{2}},\dots,C_{h_{t}}$. Unpacking definitions reveals that the distribution $\mathcal{D}_{w}$ has the property
$\Pr_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{1}{L+\alpha}$
for $t$ distinct hypotheses $h_{i}\in T$. Because these $h_{i}$ are distinct, the events $[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]$ are disjoint, and therefore
$1\geq\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}\bigcup_{i=1}^{t}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]=\sum_{i=1}^{t}\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{t}{L+\alpha}.$
It follows that $L+\alpha>t=\lceil\frac{d}{2}+1\rceil$ for any $\alpha>0$, which implies the desired lower bound $L\geq\lceil\frac{d}{2}+1\rceil$.
3.2 The Upper Bound
To prove the upper bound, we design a list replicable learning algorithm $\bm{\mathcal{A}}$ that learns $\mathcal{H}_{\gamma}^{d}$ with list size $d$ independent of $\epsilon>0$. Given $w\in\mathbb{S}^{d-1}$, let $\overline{h}_{w}:\mathbb{S}^{d-1}\to\{\pm 1\}$ denote the total concept class corresponding to the closed half-space defined by $w$.
$\overline{h}_{w}(x)\coloneqq\begin{cases}1&\text{if }\langle w,x\rangle\geq 0\\ -1&\text{if }\langle w,x\rangle<0\end{cases}.$
Fundamentally, as in [KKL+24], we estimate a large-margin linear separator using the average of many runs of an SVM maximum margin separator. Then, we use a rounding scheme based on a uniform triangulation of the $\ell_{1}$ sphere, with the guarantee that with high probability, our learning rule will choose one of at most $d$ separators.
Consider a training sample $(x_{1},y_{1}),\dots,(x_{n},y_{n})\in\mathbb{R}^{d}\times\{\pm 1\}$. The (homogeneous) hard-SVM is an optimization problem that returns a homogeneous half-space that classifies the training sample correctly while maximizing the margin $\gamma$. More formally, it is the following optimization problem over the variables $\gamma\in\mathbb{R}$ and $w\in\mathbb{S}^{d-1}$:
$\begin{array}[]{l rll}\max&\gamma&&\\ \text{s.t. }&y_{i}{\left\langle x_{i},w\right\rangle}&\geq\gamma&\text{for }i=1,\ldots,n\\ &\gamma&\geq 0&\\ &w&\in\mathbb{S}^{d-1}&\\ \end{array}$
One can use semi-definite programming to solve this optimization problem efficiently—to check whether it is feasible and, if so, to find the maximizing $w$.
Definition 3.3 ($\gamma$-Separator).
Let $S\subseteq\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$. We call $w\in\mathbb{S}^{d-1}$ a $\gamma$-separator for $S$ if
$y{\left\langle x,w\right\rangle}\geq\gamma\ \ \text{ for all }(x,y)\in S.$
Furthermore, for any distribution $\mathcal{D}$ over $\mathbb{S}^{d-1}\times\{\pm 1\}$, we call $w$ a $(\gamma,\epsilon)$-separator for $\mathcal{D}$ if
$\Pr_{(x,y)\sim\mathcal{D}}[y\langle x,w\rangle<\gamma]\leq\epsilon.$
When learning $\mathcal{H}_{\gamma}^{d}$ in the realizable setting, for any sample set $S$ drawn from a realizable distribution $\mathcal{D}$, there is some $w$ that $\gamma$-separates $S$. Therefore, the hard-SVM will be feasible and return a vector $w$ that $\gamma$-separates $S$.
The following theorem, due to [STBWA98], says that if we take a sufficiently large sample $\bm{S}$ and compute a good separator $w$ for it using hard-SVM, then with high probability, $w$ will also be a good separator for $\mathcal{D}$.
Theorem 3.4 (SVM generalization bound [STBWA98, Theorem 3.5]).
For all $\epsilon,\delta>0$, there exists $n\coloneqq n(\epsilon,\delta)$ such that the following holds. Let $\mathcal{D}$ be any distribution over $\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$.
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}\left[\text{Every $w\in\mathbb{S}^{d-1}$ that $\gamma$-separates $\bm{S}$ also $\left(\frac{\gamma}{2},\epsilon\right)$-separates $\mathcal{D}$}\right]\geq 1-\delta.$
Remark 3.5.
To prove Theorem 3.4, one can apply [STBWA98, Theorem 3.5] to show that, with probability at least $1-\delta$, both $h_{1}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}+\frac{\gamma}{2})$ and $h_{2}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}-\frac{\gamma}{2})$ have loss at most $\frac{\epsilon}{2}$, in which case $w$ is a $\left(\frac{\gamma}{2},\epsilon\right)$-separator for $\mathcal{D}$.
Regarding optimal bounds on $n(\epsilon,\delta)$ in Theorem 3.4, we refer the reader to [GKL20, KKL+24].
We will also use the following simple concentration result for sums of i.i.d. random vectors to show that the outputs of multiple runs of hard-SVM on independent samples are typically concentrated around their mean.
Lemma 3.6.
Let $\bm{x}_{1},\dots,\bm{x}_{k}\in\mathbb{R}^{d}$ be i.i.d random variables with mean $\mu$ and $\|\bm{x}_{i}-\mu\|_{\infty}\leq C$. Let $\bm{Z}=\frac{1}{k}\sum_{i=1}^{k}\bm{x}_{i}$. For all $t>0$,
$\Pr[\|\bm{Z}-\mu\|_{1}\geq t]\leq 2de^{\frac{-kt^{2}}{2d^{2}C^{2}}}.$
Proof.
We apply Hoeffding’s inequality to each coordinate and take the union bound. By Hoeffding’s inequality, for every $j\in[d]$, we have
$\Pr\left[|\bm{Z}_{j}-\mu_{j}|\geq\frac{t}{d}\right]\leq 2e^{\frac{-kt^{2}}{2d^{2}C^{2}}}.$
Therefore, by the union bound,
$\Pr\left[\|\bm{Z}-\mu\|_{1}\geq t\right]\leq 2de^{\frac{-kt^{2}}{2d^{2}C^{2}}}.\qed$
We will use a rounding scheme that ensures any small neighbourhood on $\mathbb{S}^{d-1}$ is rounded to at most $d$ points.
Lemma 3.7.
For every $\alpha>0$, there is a $\beta(d)>0$ and a rounding scheme
$\mathrm{round}_{\alpha}:\mathbb{S}^{d-1}\to\mathbb{S}^{d-1}$
such that for all $x\in\mathbb{S}^{d-1}$,
  1. 1.
    $\|\mathrm{round}_{\alpha}(x)-x\|_{2}<\alpha$, and
  2. 2.
    The set $R_{x}\coloneqq\left\{\mathrm{round}_{\alpha}(y)\ |\ y\in\mathbb{S}^{d-1}\text{and }\|x-y\|_{2}\leq\beta\right\}$ has size at most $d$.
Proof.
Consider any $\frac{\alpha}{2}$-net $T$ of points in general position on $\mathbb{S}^{d-1}$ and define the rounding function as
$\mathrm{round}_{\alpha}(x)\coloneqq\arg\min_{y\in T}\|x-y\|_{2}.$
Property 1 follows directly from the definition of $\mathrm{round}_{\alpha}$, so it remains to prove 2.
If $|T|\leq d$, both conditions are satisfied. Thus, assume $|T|>d$. We will use the fact that for any set of $d+1$ distinct points $x_{1},\ldots,x_{d+1}\in T$, the origin is the only point equidistant from all of them. To see this, suppose there exists a point $y\in\mathbb{R}^{d}$ that is equidistant from each $x_{i}$, meaning there exists some $r$ such that
$r^{2}=\|x_{i}-y\|_{2}^{2}=1+\|y\|_{2}^{2}-2{\left\langle x_{i},y\right\rangle}.$
Consequently, $y$ is orthogonal to the linearly independent vectors $x_{1}-x_{2}$, …, $x_{1}-x_{d+1}$, and thus $y=\vec{0}$.
Define the map $\phi:\mathbb{S}^{d-1}\to\mathbb{R}_{\geq 0}$ as
$\phi(x)\coloneqq\tau(x)-\min_{y\in T}\|x-y\|_{2},$
where $\tau(x)$ denotes the distance from $x$ to a $(d+1)$-th closest point in $T$. Since no point in $\mathbb{S}^{d-1}$ can be equidistant to more than $d$ points in $T$, we have $\phi(x)>0$ for all $x$. And since $\phi$ is continuous and $\mathbb{S}^{d-1}$ is compact, we have
$\beta^{\prime}\coloneqq\min_{x}\phi(x)>0.$
Taking $\beta\coloneqq\beta^{\prime}/3$ completes the proof. ∎
Upper bound of Theorem 1.6.
We need to show that for any margin $\gamma\in(0,1)$, accuracy parameter $\epsilon\in(0,1/2)$, and dimension $d\geq 1$, we have $\operatorname{LR}_{\epsilon}(\mathcal{H}_{\gamma}^{d})\leq d$.
We will construct a list-replicable learner that always outputs a hypothesis of the form $\overline{h}_{w}$ for some $w\in\mathbb{S}^{d-1}$.
Let $k=k(d,\gamma)$ and $n_{0}=n_{0}(\epsilon,\delta,k)$ be integers yet to be determined. Consider the following learning rule $\bm{\mathcal{A}}$ that uses the rounding scheme of Lemma 3.7.
Algorithm 1 The learning rule $\bm{\mathcal{A}}$
1:  for $i\leftarrow 1$ to $k$ do
2:     Sample $\mathbf{S}_{i}\sim\mathcal{D}^{n_{0}}$.
3:     Let $\bm{w}_{i}\leftarrow\text{hard-SVM}(\mathbf{S}_{i})$.
4:  end for
5:  Let $\bm{w}\leftarrow\frac{1}{k}\sum_{i=1}^{k}\bm{w}_{i}$ and $\bm{z}\leftarrow\frac{\bm{w}}{\|\bm{w}\|_{2}}$.
6:  Let $\tilde{\bm{z}}\leftarrow\mathrm{round}_{\gamma/2}(\bm{z})$.
7:  output the hypothesis $\overline{h}_{\tilde{\bm{z}}}$.
We first show that the learning rule $\bm{\mathcal{A}}$ presented in Algorithm 1 is a PAC learner.
Claim 3.8.
Let $\bm{\mathcal{A}}$ and $\bm{w}$ be as in Algorithm 1. For every $\epsilon,\delta\in(0,1)$ and $k\in\mathbb{N}$, there exists $n_{0}\coloneqq n_{0}(\epsilon,\delta,k)\in\mathbb{N}$ such that for every distribution $\mathcal{D}$ realizable by $\mathcal{H}_{\gamma}^{d}$, we have
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}\|_{2}<\frac{\gamma}{2}\right]\leq\frac{\delta}{4}$ (4)
and
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}[\mathcal{L}_{\mathcal{D}}(\bm{\mathcal{A}}(\mathbf{S}))\geq\epsilon]\leq\frac{\delta}{4}.$ (5)
Proof.
Let $\bm{w}_{1},\dots,\bm{w}_{k}$ be as in Algorithm 1. Since $\mathcal{D}$ is realizable by $\mathcal{H}_{\gamma}^{d}$, for every $i$, $\bm{w}_{i}$ is a $\gamma$-separator for $\mathbf{S}_{i}$. Therefore, by Theorem 3.4, if $n_{0}(\epsilon,\delta,k)$ is sufficiently large,
$\Pr_{\begin{subarray}{c}\mathbf{S}_{i}\sim\mathcal{D}^{n_{0}}\end{subarray}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\right]\geq 1-\frac{\delta}{4k}.$
Thus, by the union bound,
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\text{ for all $i\in[k]$}\right]\geq 1-\frac{\delta}{4},$ (6)
and applying the union bound again,
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\min_{i\in[k]}\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\epsilon\right]\geq 1-\frac{\delta}{4}.$ (7)
Finally, if $(x,y)\in\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$ satisfies $y{\left\langle x,\bm{w}_{i}\right\rangle}\geq\gamma/2$ for all $i\in[k]$, then noting that $\|\bm{w}\|\leq 1$, we have
$y{\left\langle x,\bm{z}\right\rangle}\geq y{\left\langle x,\bm{w}\right\rangle}=y{\left\langle x,\frac{1}{k}\sum_{i=1}^{k}\bm{w}_{i}\right\rangle}\geq\gamma/2.$ (8)
Thus, from Equation 7, we have
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}\|_{2}<\frac{\gamma}{2}\right]\leq\frac{\delta}{4}$
and
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{z}\right\rangle}\geq\frac{\gamma}{2}\right]\geq 1-\epsilon\right]\geq 1-\frac{\delta}{4}.$ (9)
By applying Lemma 3.7 with $\alpha\coloneqq\gamma/2$, after rounding $\bm{z}$ to $\tilde{\bm{z}}$, we have $\|\tilde{\bm{z}}-\bm{z}\|_{2}<\frac{\gamma}{2}$. Thus if $(x,y)\in\mathbb{S}^{d-1}\times\{\pm 1\}$ satisfy $y{\left\langle x,\bm{z}\right\rangle}\geq\gamma/2$, then
$y{\left\langle x,\tilde{\bm{z}}\right\rangle}=y{\left\langle x,\bm{z}\right\rangle}+y{\left\langle x,\tilde{\bm{z}}-\bm{z}\right\rangle}\geq\frac{\gamma}{2}-\|\tilde{\bm{z}}-\bm{z}\|_{2}>0,$
namely $\overline{h}_{\tilde{\bm{z}}}(x)=y$. Thus,
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\mathcal{L}_{\mathcal{D}}(\overline{h}_{\tilde{\bm{z}}})\leq\epsilon\right]\geq 1-\frac{\delta}{4},$
which completes the proof of the claim. ∎
We now complete the proof by addressing list replicability. Let $\beta$ be as in Lemma 3.7. Applying Lemma 3.6, since $\bm{w}$ is the average of $k$ i.i.d. random variables in $\mathbb{S}^{d-1}$, there exists $k=k(\gamma,d)\in\mathbb{N}$ such that
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}-\operatorname{\mathbb{E}}[\bm{w}]\|_{2}\geq\frac{\gamma\beta}{2}\right]\leq\frac{\delta}{4}.$
Since $\bm{z}=\frac{\bm{w}}{\|\bm{w}\|_{2}}$, by applying the union bound to Eq. 4 and the above inequality, we have
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{z}-\operatorname{\mathbb{E}}[\bm{z}]\|_{2}\geq\beta\right]\leq\frac{\delta}{2}.$
Consequently, by Lemma 3.7, with probability at least $21-\delta/$ 2, the rounding scheme ($\mathrm{round}_{\frac{\gamma}{2}}$) outputs one of at most $d$ hypotheses. Applying a union bound with 3.8 completes the proof of the upper bound of Theorem 1.6. ∎
4 Disambiguations of Gap Hamming Distance
We first prove that every disambiguation of the Gap Hamming Distance matrix family has unbounded public-coin randomized communication complexity.
Proof of Corollary 1.15.
We combine Theorem 1.14 with the known relationship between Littlsetone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity. Given a matrix $M\in\{\pm 1\}^{\mathcal{X}\times\mathcal{Y}}$, the margin of $M$ is defined
$\operatorname{m}(M)\coloneqq\max_{\begin{subarray}{c}d\in\mathbb{N},\\ u_{x},u_{y}\in\mathbb{S}^{d}\end{subarray}}\min_{(x,y)}M(x,y)\cdot{\left\langle u_{x},u_{y}\right\rangle}.$
In other words, $\operatorname{m}(M)$ is the largest $\gamma$ such that $M$ appears as a submatrix of $\mathcal{H}^{d}_{\gamma}$ for some $d$.
Let $\{M_{n}\}_{n=1}^{\infty}$ be a family of disambiguation of $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$. By Theorem 1.14, we know that $\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n})$. It thus follows from LABEL:{Ldim_of_margin} that
$\operatorname{m}(M_{n})=O\left(\frac{1}{\sqrt[4]{\log n}}\right).$
Finally, invoking the equivalence of margin and discrepancy by [LS09] and the relation between discrepancy and randomized communication complexity by [CG88] (see also [HHP+22, Proposition 3.3 ]) shows that the public-coin randomized communication complexity of $M_{n}$ is
$\Omega(\log(\operatorname{m}(M_{n})^{-1}))=\Omega(\log\log n).\qed$
We next present the proof of Theorem 1.14. The key is to use an embedding of bounded margin half-spaces in dimension $d$ into the Boolean cube, which allows us to disambiguate $\mathcal{H}^{d}_{\gamma}$ using a disambiguation of the Gap Hamming Distance problem in dimension $O(d)$. The existence of such an embedding was proved in [HHM23], which we rephrase as follows.
Lemma 4.1 (Adapted from [HHM23, Lemma 3.2]).
Let $\gamma\in(0,1)$ and $n\in\mathbb{N}$. There exist $d=\Omega\left((1-\gamma)^{2}\cdot n/\log(1/(1-\gamma))\right)$, $\gamma^{\prime}\in(0,1)$ and a map $\xi\colon\mathbb{S}^{d-1}\to\left\{\pm 1\right\}^{n}$ such that for all $u,v\in\mathbb{S}^{d-1}$, we have
$\langle u,v\rangle\geq\gamma^{\prime}\implies\langle\xi(u),\xi(v)\rangle\geq\gamma n$
$\langle u,v\rangle\leq-\gamma^{\prime}\implies\langle\xi(u),\xi(v)\rangle\leq-\gamma n$
We are now ready to prove Theorem 1.14.
Proof of Theorem 1.14.
Fix $\gamma\in(0,1)$. Let $\{M_{n}\}_{n=1}^{\infty}$ be a family of total functions which disambiguates $\{\operatorname{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$, and let $d=d(n)$ and $\gamma^{\prime}$ be as provided by Lemma 4.1.
We will use this lemma along with the functions $\left\{M_{n}\right\}_{n=1}^{\infty}$ to disambiguate the family of partial concept classes $\{\mathcal{H}_{\gamma^{\prime}}^{d(n)}\}_{n=1}^{\infty}$. To this end, we disambiguate each partial concept $h_{w}\in\mathcal{H}_{\gamma^{\prime}}^{d}$ (defined in LABEL:{eq:def_large_margin}) to
$\overline{h}_{w}(x)\coloneqq M_{n}\big{(}\xi(w),\xi(x)\big{)}.$
Let us verify that $\overline{h}_{w}$ is, in fact, a disambiguation of $h_{w}$.
Suppose that $h_{w}(x)=1$. By definition, this occurs exactly when $\langle w,x\rangle\geq\gamma^{\prime}$. It follows from the properties of $\xi$ that
$\langle\xi(w),\xi(x)\rangle\geq\gamma n.$
Therefore, for such $w,x$, we have
$\overline{h}_{w}(x)=M_{n}(\xi(w),\xi(x))=\operatorname{GHD}_{\gamma}^{n}(\xi(w),\xi(x))=1=h_{w}(x).$
A similar argument shows that if $h_{w}(x)=-1$, then $\overline{h}_{w}(x)=-1$. We deduce that $\overline{h}_{w}$ indeed disambiguates $h_{w}$, and ¯$\overline{\mathcal{H}}^{d}_{\gamma^{\prime}}$ is a disambiguation of $\mathcal{H}^{d}_{\gamma^{\prime}}$.
Finally, note that by construction, any shattered mistake tree in ¯$\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}$ corresponds to a shattered mistake tree of the same depth in $M_{n}$. Therefore, $\operatorname{Ldim}(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d})\leq\operatorname{Ldim}(M_{n})$. This combined with Theorem 1.8 implies that
$\operatorname{Ldim}(M_{n})\geq\operatorname{Ldim}\left(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}\right)=\Omega\left(\sqrt{\log d(n)}\right)=\Omega\left(\sqrt{\log n}\right).\qed$
Appendix A Equivalence of Stability and List-Replicability
We prove the following more refined version of Theorem 1.2.
Theorem A.1.
Let $\mathcal{H}$ be any total or partial concept class on the domain $\mathcal{X}$. Then for every $\epsilon\in(0,1)$,
$\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}\qquad\text{ and }\qquad\operatorname{LR}_{\epsilon}(\mathcal{H})\leq\frac{1}{\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{H})}.$
Consequently, $\rho^{\mathrm{gs}}(\mathcal{H})=\frac{1}{\operatorname{LR}(\mathcal{H})}$.
Proof.
We first prove that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}$. Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon,L)$-list replicable learner for $\mathcal{H}$ with sample complexity $n=n(\epsilon,\delta)$. Let $\mathcal{D}$ be any realizable distribution on $\mathcal{X}\times\{\pm 1\}$, and let $h_{1},\ldots,h_{L}$ be the list of hypotheses guaranteed by Definition 1.1.
By the pigeonhole principle, at least one $h_{i}$ satisfies
$\Pr_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})=h_{i}]\geq\frac{1-\delta}{L}.$
Since this statement holds for arbitrary $\delta>0$, $\bm{\mathcal{A}}$ is itself an $(\epsilon,\rho)$-globally stable learner for all $\rho<\frac{1}{L}$. We may conclude that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}$.
Next, we prove that $\operatorname{LR}_{\epsilon}(\mathcal{H})\leq 1/\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{H})$. Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon/3,\rho)$-globally stable learner for $\mathcal{H}$ with sample complexity $n_{0}=n_{0}(\epsilon)$. By the stability assumption, for every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\{\pm 1\}$, there exists $h^{*}:\mathcal{X}\to\left\{\pm 1\right\}$ satisfying
$\mathcal{L}_{\mathcal{D}}(h^{*})\leq\frac{\epsilon}{3}\ \text{ and }\ \Pr_{\begin{subarray}{c}\bm{S}\sim\mathcal{D}^{n_{0}}\end{subarray}}[\bm{\mathcal{A}}(\bm{S})=h^{*}]\geq\rho.$ (10)
For every $h\in\{\pm 1\}^{\mathcal{X}}$ and realizable distribution $\mathcal{D}$, define
$p(h)\coloneqq\Pr_{\mathbf{S}\sim\mathcal{D}^{n_{0}}}[\bm{\mathcal{A}}(\bm{S})=h],$
Denote $L\coloneqq\left\lfloor\frac{1}{\rho}\right\rfloor$, so that $\rho\in\left(\frac{1}{L+1},\frac{1}{L}\right]$, and let $\alpha\coloneqq\rho-\frac{1}{L+1}>0$. Define the list $\Lambda$ of good and likely hypotheses
$\Lambda\coloneqq\left\{h\in\left\{\pm 1\right\}^{\mathcal{X}}\ |\ p(h)>\frac{1}{L+1}\text{ and }\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right\}.$
Note that $|\Lambda|\leq L$ and $\Lambda$ is nonempty, as it contains $h^{*}$. Therefore, to construct an $(\epsilon,L)$-list replicable learner, it suffices to show that for any confidence parameter $\delta>0$, the learning rule outputs a hypothesis from $\Lambda$ with probability at least $1-\delta$.
Let $t\coloneqq t(\alpha,\delta)$ and $n_{1}\coloneqq n_{1}(\epsilon,t)$ be sufficiently large integers to be determined later. We propose the following learning rule $\bm{\mathcal{A}}^{\prime}$ with sample complexity $tn_{0}+n_{1}$.
Algorithm 2 The learning rule $\bm{\mathcal{A}}^{\prime}$
1:  Sample a dataset:
$\bm{S}=({\bm{P}},\bm{Q})\sim\mathcal{D}^{tn_{0}+n_{1}},\quad\text{where }{\bm{P}}=({\bm{P}}_{1},\ldots,{\bm{P}}_{t})\sim(\mathcal{D}^{n_{0}})^{t}=\mathcal{D}^{tn_{0}},\quad\text{and}\quad\bm{Q}\sim\mathcal{D}^{n_{1}}.$
2:  Define the empirical estimate of $p(h)$ as
$\operatorname{freq}_{\bm{P}}(h)\coloneqq\frac{|\left\{i\in[t]\ |\ \bm{\mathcal{A}}({\bm{P}}_{i})=h\right\}|}{t}.$
3:  Output any hypothesis $h\in\{\pm 1\}^{\mathcal{X}}$ satisfying:
  • $\operatorname{freq}_{\bm{P}}(h)\geq\rho-\tfrac{\alpha}{2}$
  • $\mathcal{L}_{{\bm{Q}}}(h)\leq\tfrac{2\epsilon}{3}$
If no such $h$ exists, output an arbitrary $h$ corresponding to “failure.”
Denote by $\mathcal{Y}$ the set of all $h$ with $\operatorname{freq}_{\bm{P}}(h)>0$ in Algorithm 2, and note that $|\mathcal{Y}|\leq t$. To show that $\bm{\mathcal{A}}^{\prime}$ outputs a hypothesis from $\Lambda$ with probability at least $1-\delta$, we will condition on the events
$A:\quad$ $|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}$
$B:\quad$ $|p(h)-\operatorname{freq}_{\bm{P}}(h)|<\frac{\alpha}{2}\text{ forall }h\in\left\{\pm 1\right\}^{\mathcal{X}}$
To guarantee that both events are likely, we prove the following claim.
Claim A.2.
There exist integers $t(\alpha,\delta)$ and $n_{1}(\epsilon,t)$ such that
$\Pr_{{\bm{P}}\sim\mathcal{D}^{tn_{0}}}[B]\geq 1-\frac{\delta}{2}\quad\text{and}\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}[A]\geq 1-\frac{\delta}{2}.$
Proof of A.2.
For the choice of $t$ and the proof of the first inequality, we use the uniform convergence property of the family of indicator functions on {±$\{\pm 1\}^{\mathcal{X}}$. More precisely, for $f\in\left\{\pm 1\right\}^{\mathcal{X}}$, define $\mathbb{I}_{f}:\left\{\pm 1\right\}^{\mathcal{X}}\to\left\{0,1\right\}$ as
$\mathbb{I}_{f}(f^{\prime})\coloneqq\begin{cases}1&f^{\prime}=f\\ 0&\text{otherwise}\end{cases}.$
The class
$\mathcal{I}\coloneqq\left\{\mathbb{I}_{f}\ |\ f\in\left\{\pm 1\right\}^{\mathcal{X}}\right\}$
has VC dimension $1$, and therefore, it satisfies the uniform convergence property. For ${\bm{P}}_{i}\sim\mathcal{D}^{n_{0}}$, $\bm{\mathcal{A}}({\bm{P}}_{i})$ induces a probability distribution $\mu$ on {±$\left\{\pm 1\right\}^{\mathcal{X}}$, and we have
$1-p(h)=\Pr_{{\bm{P}}_{i}\sim\mathcal{D}^{n_{0}}}[\bm{\mathcal{A}}({\bm{P}}_{i})\neq h]=\mathcal{L}_{\mu}(\mathbb{I}_{h}),$
while $1-\operatorname{freq}_{\bm{P}}(h)$ corresponds to the empirical loss of $\mathbb{I}_{h}$ on $(\mathbb{I}_{h_{1}},\ldots,\mathbb{I}_{h_{t}})\sim\mu^{t}$. Thus, by the uniform convergence property on $\mathcal{I}$, our claim holds.
Now that we have $t$, we can define $n_{1}$ and prove the second inequality. Note that for every $h\in\{\pm 1\}^{\mathcal{X}}$, for ${\bm{Q}}\sim\mathcal{D}^{n_{1}}$, $\mathcal{L}_{\bm{Q}}(h)$ is an average of ${n_{1}}$ samplings of a Bernoulli random variable with expectation $\mathcal{L}_{\mathcal{D}}(h)$. Thus, by Hoeffding’s inequality, there exists $n_{1}=n_{1}(\epsilon^{\prime},t)$ such that
$\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}\left[\left|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)\right|>\frac{\epsilon}{3}\right]\leq\frac{\delta}{2t}.$ (11)
Thus, by the union bound, we have
$\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}\left[\left|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)\right|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}\right]\geq 1-\frac{\delta}{2}.$ (12)
A direct consequence of Claim A.2 is that
$\Pr_{\mathbf{S}\sim\mathcal{D}^{tn_{0}+n_{1}}}\left[A,B\right]\geq 1-\delta.$
Condition on events $A$ and $B$, and let $h^{*}$ be a stable hypothesis for $\bm{\mathcal{A}}$, as described in Eq. 10. We will show that $h^{*}$ is a candidate for output, so $\bm{\mathcal{A}}^{\prime}$ will not output “failure”. To check the first condition for output, we combine $B$ and Eq. 10 to show that
$\operatorname{freq}_{\bm{P}}(h^{*})\geq p(h^{*})-\frac{\alpha}{2}\geq\rho-\frac{\alpha}{2}.$
Moreover, $\rho-\frac{\alpha}{2}>0$, so $h^{*}\in\mathcal{Y}$. We may therefore apply $A$ to show that $h^{*}$ satisfies the second condition for output,
$\mathcal{L}_{\bm{Q}}(h^{*})\leq\mathcal{L}_{\mathcal{D}}(h^{*})+\frac{\epsilon}{3}\leq\frac{2\epsilon}{3}.$
Finally, let $h_{o}$ be any output of $\bm{\mathcal{A}}^{\prime}$, conditioned on $A$ and $B$. Then, $h_{o}$ satisfies the condition $\operatorname{freq}_{\bm{P}}(h_{o})\geq\rho-\frac{\alpha}{2}$, so because of $B$,
$p(h_{o})>\operatorname{freq}_{\bm{P}}(h_{o})-\frac{\alpha}{2}\geq\rho-\alpha=\frac{1}{L+1}.$
Furthermore, $h_{o}$ also satisfies the condition $\mathcal{L}_{\bm{Q}}(h_{o})\leq\frac{2\epsilon}{3}$, so because of $A$,
$\mathcal{L}_{\mathcal{D}}(h_{o})\leq\mathcal{L}_{\bm{Q}}(h_{o})+\frac{\epsilon}{3}\leq\epsilon.$
Thus, $h_{o}$ must be in $\Lambda$. ∎

### v4

@@
 Abstract
-Recent remarkable advances in learning theory have established that, for total concept classes, list replicability, global stability, differentially private (DP) learnability, and shared-randomness replicability all coincide with the finiteness of Littlestone dimension.
-Does this equivalence extend to partial concept classes?
-We answer this question by proving that the list replicability number of $d$-dimensional $\gamma$-margin half-spaces satisfies
+We prove that the list replicability number of $d$-dimensional $\gamma$-margin half-spaces satisfies
 $\frac{d}{2}+1\leq\operatorname{LR}(\mathcal{H}^{d}_{\gamma})\leq d,$
-which grows with dimension.
-Consequently, for partial classes, list replicability and global stability do not necessarily follow from bounded Littlestone dimension, pure DP-learnability, or shared-randomness replicability.
-Applying our main theorem, we resolve several open problems:
+which grows with dimension. This resolves several open problems:
   • Every disambiguation of infinite-dimensional large-margin half-spaces to a total concept class has unbounded Littlestone dimension, answering an open question of Alon, Hanneke, Holzman, and Moran (FOCS ’21).
-  • The maximum list-replicability number of any finite set of points and homogeneous half-spaces in $d$-dimensional Euclidean space is $d$, resolving a problem of Chase, Moran, and Yehudayoff (FOCS ’23).
   • Every disambiguation of the Gap Hamming Distance problem in the large gap regime has unbounded public-coin randomized communication complexity.
 This answers an open question of Fang, Göös, Harms, and Hatami (STOC ’25).
+  • There is a separation of $O(1)$ vs $\omega(1)$ between randomized and pseudo-deterministic communication complexity.
+  • The maximum list-replicability number of any finite set of points and homogeneous half-spaces in $d$-dimensional Euclidean space is $d$, resolving a problem of Chase, Moran, and Yehudayoff (FOCS ’23).
   • There exists a partial concept class with Littlestone dimension $1$ such that all its disambiguations have infinite Littlestone dimension.
-This answers a question of Cheung, H.
+This resolves a problem of Cheung, H.
 Hatami, P.
 Hatami, and Hosseini (ICALP ’23).
-Our lower bound follows from a topological argument based on the local Borsuk-Ulam theorem of Chase, Chornomaz, Moran, and Yehudayoff (STOC ’24).
-For the upper bound, we construct a list-replicable learning rule using the generalization properties of SVMs.
+Our lower bound follows from a topological argument based on a local Borsuk-Ulam theorem. For the upper bound, we construct a list-replicable learning rule using the generalization properties of SVMs.
 1 Introduction
-The large-margin half-space classification problem is a central topic in learning theory, extensively studied for its theoretical significance and practical impact.
-Theoretically, it is a fundamental and elegant geometric problem, and it serves as a simple model for understanding more complex learning problems.
-Practically, its importance stems from the success of Support Vector Machines (SVMs), which leverage the large-margin assumption to produce accurate classifications in high-dimensional spaces.
-SVMs play a key role in many real-world applications, including pattern and text recognition, image analysis, geographic information systems, bioinformatics, healthcare, and fraud detection.
-To describe the large-margin classification problem, we first consider standard half-space classification without any margin assumption.
-After appropriate normalization and homogenization, we may assume that the domain is the unit sphere $\mathbb{S}^{d-1}\subseteq\mathbb{R}^{d}$, and the objective is to learn an unknown homogeneous half-space which classifies each point on the sphere as either ±$\pm 1$ depending on whether it belongs to the half-space.
+Large-margin half-space classification is a fundamental problem in learning theory.
+In this setting, data is normalized to lie on the unit sphere $\mathbb{S}^{d-1}\subset\mathbb{R}^{d}$, and we are guaranteed a promise that each point lies at least a fixed margin $\gamma\in(0,1)$ from some unknown homogeneous hyperplane.
+The learner is then tasked with classifying points based on the side of the defining hyperplane to which they belong.
+This problem has been extensively studied for both its theoretical and practical significance: it provides a clean geometric model for analyzing more complex learning tasks, and underlies the success of Support Vector Machines (SVMs), which leverage the large-margin assumption to produce accurate classifications in high-dimensional spaces, with applications across domains such as text and image recognition, bioinformatics, and fraud detection.
+We study this problem through the lens of replicability, the requirement that an algorithm produce consistent outcomes when repeated under similar conditions and with similar data.
+In recent years, replicability has become a vibrant research area, and various rigorous formulations of replicability for learning algorithms have been introduced and studied [BLM20, MM22, CMY23, BGH+23, KVYZ23, EKK+23, EKM+23, MSS23, EHKS23, KKL+24, KKMV23].
+Among them, one of the most intriguing is the notion of global stability, which was first discovered in connection with differentially private learning and online learning [BLM20, ABL+22, CMY23].
+Subsequent work, however, has shown that its significance extends well beyond these applications.
+In its equivalent formulation as list replicability [CMY23], the notion is intrinsically linked to the geometry and topology of the space of realizable distributions of a concept class [CMY23, CCMY24, BGHH25, CMW25, BHH+25].
+For example, for finite classes, it characterizes the topological dimension of this space under its natural simplicial structure [BHH+25].
+Through these connections, fundamental results in classical topological dimension theory, such as the Lebesgue covering theorem, translate directly into statements about learnability and replicability.
+Our main result states that the list replicability number of the large-margin classification problem in $\mathbb{R}^{d}$ lies between $\tfrac{d}{2}+1$ and $d$.
+In particular, it diverges as the ambient dimension $d$ grows.
+This stands in contrast to many common complexity measures for the same task, such as the VC dimension, Littlestone dimension, and randomized communication complexity, which are bounded by a function of the margin $\gamma\in(0,1)$ independent of the ambient dimension $d$.
+This divergence has several consequences and resolves a number of open problems from previous works, as discussed in detail in Section 2.1.
+Beyond its implications within learning theory, it also connects naturally to questions in communication complexity.
+In particular, our most notable consequence shows that any disambiguation of the large-margin classification problem into a total concept class must have a large Littlestone dimension and a large randomized communication complexity.
+Thus, while the original partial problem is “easy” under these classical measures, every possible completion of it to a total problem is inherently “hard”.
+Establishing such lower bounds for disambiguations is typically very challenging, as the initial partial problem is “easy” and one has no control over how it is extended to a total one.
+A notable consequence of our disambiguation theorem is an $O(1)$ versus $\Omega(\log\log n)$ gap between randomized and pseudo-deterministic communication complexities.
+Separating these two measures is a well-known open problem in the study of pseudodeterminism [GIPS21], and our result provides the first $O(1)$ versus $\omega(1)$ separation.
+1.1 Preliminaries
+We study the large-margin half-space problem through the formal lens of partial classes, which offers a general framework for analyzing such constrained learning tasks.
+A partial concept class over an arbitrary domain $\mathcal{X}$ is a set $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, where each $c\in\mathcal{C}$ is called a partial concept.
+The value $c(x)=\star$ indicates that $c$ is undefined at $x$, and therefore, both $\pm 1$ are acceptable predictions for the label of $x$.
+PAC learning.
 The standard mathematical framework for analyzing the complexity of a learning task is probably approximately correct (PAC) learning.
-In this setting, the learner receives a training set of labeled points drawn independently from an unknown distribution $\mathcal{D}$, and the goal is to produce a hypothesis that accurately predicts the labels of new points sampled from $\mathcal{D}$ with high probability.
-The fundamental theorem of PAC learning states that the size of the training set required for PAC learning depends on a combinatorial parameter known as the VC dimension (Definition 2.5).
-We also consider the Littlestone dimension (Definition 2.7), a refinement of the VC dimension that determines the optimal number of mistakes in online learning.
-In online learning, the learner receives data points sequentially and must predict the label of each point in turn before observing its correct label.
-The goal is to minimize the number of mistakes throughout the learning process.
-Since the VC dimension of homogeneous half-spaces in $\mathbb{R}^{d}$ is $d$, PAC-learning this class requires a large training set in the high-dimensional settings.
-Moreover, the Littlestone dimension of this class is infinite except in the trivial case of $d=1$, meaning the class is not online learnable, even in $\mathbb{R}^{2}$.
-Under the large-margin assumption, the learning task is restricted to points that lie at least a margin $\gamma>0$ away from the defining hyperplane.
-This constraint allows algorithms such as Perceptron [MP43, Ros58] and Support Vector Machines [VC64, CV95] to achieve efficient PAC and online learning, even in high-dimensional spaces.
-We study the large-margin halfspaces problem through the formal lens of partial classes, which offers a general framework for analyzing such constrained learning tasks.
-A partial concept class over an arbitrary domain $\mathcal{X}$ is a set {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, where each $h\in\mathcal{H}$ is called a partial concept.
-The value $h(x)=\star$ indicates that $h$ is undefined at $x$, and therefore, both ±$\pm 1$ are acceptable predictions for the label of $x$.
-In the large-margin setting, the domain is $\mathbb{S}^{d-1}$ and every homogeneous half-space defines a partial concept that assigns $h(x)=\star$ if $x$ lies within distance $\gamma$ of the defining hyperplane of $h$.
-Otherwise, it classifies $x$ as ±$\pm 1$ depending on whether it belongs to the half-space.
-More formally, each hypothesis $h_{w}:\mathbb{S}^{d-1}\to\{\pm 1,\star\}$ is specified by a unit vector $w\in\mathbb{S}^{d-1}$ and given by
-$h_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}(\langle w,x\rangle)&\text{if }|\langle w,x\rangle|\geq\gamma\\ \star&\text{otherwise}\end{cases}.$ (1)
-We denote the partial concept class of all such $h_{w}$ by $\mathcal{H}^{d}_{\gamma}$.
-We also define $\mathcal{H}^{\infty}_{\gamma}$ as the class of all partial concepts $h_{w}:\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}\to\left\{\pm 1,\star\right\}$, where each hypothesis is specified by a unit vector $w\in\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}$ of arbitrary finite dimension.
-For any $x$ and $w$ of different dimensions, we set $h_{w}(x)\coloneqq\star$, and otherwise, define $h_{w}(x)$ as in Eq.
-1.
-In [AHHM22], Alon, Hanneke, Holzman, and Moran proved that, as in the case of total concept classes, the VC and Littlestone dimensions characterize PAC and online learnability of partial concept classes.
-In the case of $\mathcal{H}_{\gamma}^{d}$, the classic mistake-bound analysis of the Perceptron algorithm [MP43, Ros58] (see also [SSBD14, Theorem 9.1]) shows the following upper bound on the Littlestone and VC dimensions:
+In PAC learning, the learner is given parameters $\delta,\epsilon>0$ and receives training data consisting of $n=n(\mathcal{C},\delta,\epsilon)$ independent labeled examples drawn from an unknown but fixed distribution $\mu$ over $\mathcal{X}\times\{\pm 1\}$.
+We work in the realizable setting: for every $n$, a random sample $\bm{S}=((\bm{x}_{i},\bm{y}_{i}))_{i=1}^{n}\sim\mu^{n}$ is almost surely realizable by some $c\in\mathcal{C}$, meaning that $c(\bm{x}_{i})=\bm{y}_{i}$ for all $i=1,\ldots,n$.
+[^1] Here, and throughout the paper, we use boldface letters to denote random variables and use the notation $(\bm{x},\bm{y})\sim\mu$ to express that $(\bm{x},\bm{y})$ is a random variable distributed according to $\mu$.
+Note that $\mu$ is a distribution over $\mathcal{X}\times\left\{\pm 1\right\}$, so none of the labels $\bm{y}_{i}$ take the value $\star$.
+The learner’s task is to use the training data to output, with probability at least $1-\delta$, a hypothesis $h:\mathcal{X}\to\{\pm 1\}$ whose population loss
+$\operatorname{loss}_{\mu}(h)\coloneqq\Pr_{(\bm{x},\bm{y})\sim\mu}[h(\bm{x})\neq\bm{y}]$
+is at most $\epsilon$.
+The following simple lemma from [AHHM21] establishes the connection between realizability and having zero population loss.
+Lemma 1.1 ([AHHM21]).
+Let $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class, and let $\mu$ be a distribution on $\mathcal{X}\times\left\{\pm 1\right\}$.
+If $\operatorname{loss}_{\mu}(\mathcal{C})\coloneqq\inf_{c\in\mathcal{C}}\operatorname{loss}_{\mu}(c)$ is zero, then $\mu$ is realizable by $\mathcal{C}$.
+Conversely, if $\mu$ is realizable and has finite or countable support, then $\operatorname{loss}_{\mu}(\mathcal{C})=0$.
+The fundamental theorem of PAC learning states that the size of the training set required for PAC learning a total concept class depends on a combinatorial parameter known as the VC dimension, which we now define in the more general partial setting.
+A (partial) concept class $\mathcal{C}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$ shatters a set $S\subseteq\mathcal{X}$ if $\{c|_{S}:c\in\mathcal{C}\}=\{\pm 1\}^{S}$, where $c|_{S}$ denotes the restriction of $c$ to $S$.
+The VC dimension of $\mathcal{C}$ is defined as
+$\operatorname{VCdim}(\mathcal{C})\coloneqq\sup\{|S|:\ S\subseteq\mathcal{X}\text{ is shattered by }\mathcal{C}\}.$
+In [AHHM21], Alon, Hanneke, Holzman, and Moran proved that the fundamental theorem of PAC learning holds for partial concept classes as well.
+List replicability.
+Throughout this work, a learning rule refers to a (randomized) function $\bm{\mathcal{A}}$ that maps any sample $S\in\bigcup_{n=0}^{\infty}(\mathcal{X}\times\left\{\pm 1\right\})^{n}$ to a hypothesis $\bm{\mathcal{A}}(S)\in\left\{\pm 1\right\}^{\mathcal{X}}$.
+Since our primary focus is sample complexity rather than computational efficiency, we impose no computability constraints on $\bm{\mathcal{A}}$.
+Definition 1.2 (List replicability).
+A learning rule $\bm{\mathcal{A}}$ is an $(\epsilon,L)$-list replicable learner for $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ if for every $\delta>0$, there is a sample complexity $n=n(\delta)$ such that the following holds.
+For every realizable distribution $\mu$ on $\mathcal{X}\times\left\{\pm 1\right\}$, there exists $h_{1},\ldots,h_{L}\in\left\{\pm 1\right\}^{\mathcal{X}}$ such that
+$\operatorname{loss}_{\mu}(h_{i})\leq\epsilon~\forall i~\text{ and }~\Pr_{\bm{S}\sim\mu^{n}}[\bm{\mathcal{A}}(\bm{S})\not\in\left\{h_{1},\ldots,h_{L}\right\}]\leq\delta.$
+The $\epsilon$-list replicability number of $\mathcal{C}$ is
+$\operatorname{LR}_{\epsilon}(\mathcal{C})\coloneqq\min\{L:\exists(\epsilon,L)\text{-list replicable learner for }\mathcal{H}\},$
+with $\operatorname{LR}_{\epsilon}(\mathcal{C})=\infty$ if none exists. The list replicability number of $\mathcal{C}$ is
+$\operatorname{LR}(\mathcal{C})\coloneqq\sup_{\epsilon>0}\operatorname{LR}_{\epsilon}(\mathcal{C}).$
+We say $\mathcal{C}$ is list replicable if $\operatorname{LR}(\mathcal{C})<\infty$.
+Definition 1.2 provides a strong notion of replicability as the learner’s output is typically chosen from a small list $\left\{h_{1},\ldots,h_{L}\right\}$, and all these hypotheses have small population loss.
+Remark 1.3.
+Some readers might be familiar with an equivalent form of list replicability known as global stability.
+Its definition is not needed for the main results of this paper, so we include it in Section A.3, along with the definition of the global stability parameter $\rho(\mathcal{C})$ of a concept class $\mathcal{C}$.
+This parameter is analogous to the list replicability number, and in fact [CMY23] proved that the two quantities are related by the equation $\rho(\mathcal{C})=1/\operatorname{LR}(\mathcal{C})$ for total concept classes.
+It is easy to check that the proof extends to the partial setting (see Appendix B).
+Hence, qualitative results about list replicability also hold for global stability, and quantitative results hold after the appropriate reciprocal modification.
+Online learning and Littlestone dimension.
+In online learning, a learner receives data points sequentially from an adversary and must predict each label before seeing the correct answer.
+The goal is to minimize the total number of mistakes.
+The optimal mistake bound is captured by the Littlestone dimension, a refinement of the VC dimension defined via mistake trees.
+A mistake tree of depth $d$ over domain $\mathcal{X}$ is a complete binary tree whose internal nodes are labeled by points $x\in\mathcal{X}$ and edges by bits $b\in\{\pm 1\}$ (1-1 for left, +1+1 for right).
+Following a path from the root to a leaf thus yields a sequence $(x_{1},b_{1}),\dots,(x_{d},b_{d})$, where each $x_{i}$ is the node label at level $i$ and $b_{i}$ records whether the path goes left or right.
+A concept class $\mathcal{C}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$ shatters such a tree if for every root-to-leaf path there exists $c\in\mathcal{C}$ with $c(x_{i})=b_{i}$ for all $i$.
+The Littlestone dimension $\operatorname{Ldim}(\mathcal{C})$ is the largest $d$ for which some depth-$d$ mistake tree is shattered, or $\infty$ if no such $d$ exists.
+It always holds that $\operatorname{VCdim}(\mathcal{C})\leq\operatorname{Ldim}(\mathcal{C})$, since any set $S=\left\{x_{1},\ldots,x_{d}\right\}$ shattered by $\mathcal{C}$ gives rise to a mistake tree of depth $d$, where all nodes at level $i$ are labeled with $x_{i}$.
+This tree is shattered by $\mathcal{C}$.
+Littlestone proved that a total concept class $\mathcal{C}$ is online learnable if and only if $\operatorname{Ldim}(\mathcal{C})<\infty$.
+This result was later extended to partial concept classes by [AHHM21]
+Large-margin half-spaces.
+In the large-margin setting, the domain is $\mathbb{S}^{d-1}$ and every homogeneous half-space defines a partial concept that assigns $c(x)=\star$ if $x$ lies within distance $\gamma$ of the defining hyperplane of $h$.
+Otherwise, it classifies $x$ as $\pm 1$ depending on whether it belongs to the half-space.
+More formally, each concept $c_{w}:\mathbb{S}^{d-1}\to\{\pm 1,\star\}$ is specified by a unit vector $w\in\mathbb{S}^{d-1}$ and given by
+$c_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}({\left\langle w,x\right\rangle})&\text{if }|{\left\langle w,x\right\rangle}|\geq\gamma\\ \star&\text{otherwise}\end{cases}.$ (1)
+We denote the partial concept class of all such $c_{w}$ by $\mathcal{H}^{d}_{\gamma}$.
+For the standard half-space classification problem $\mathcal{H}^{d}$ without any margin assumption (that is, each $x$ is labeled by $\operatorname{sgn}({\left\langle w,x\right\rangle})$ whenever ${\left\langle w,x\right\rangle}\neq 0$ and by $\star$ otherwise), we have $\operatorname{VCdim}(\mathcal{H}^{d})=d$.
+Moreover, $\operatorname{Ldim}(\mathcal{H}^{d})=\infty$, except in the trivial case of $d=1$.
+In particular, this class is not online learnable, even in $\mathbb{R}^{2}$.
+However, under the large-margin assumption $\gamma>0$, the classic mistake-bound analysis of the Perceptron algorithm [MP43, Ros58] (see also [SSBD14, Theorem 9.1]) shows the following upper bound on the Littlestone and VC dimensions:
 $\operatorname{VCdim}(\mathcal{H}_{\gamma}^{d})\leq\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})\leq\gamma^{-2}.$ (2)
 Crucially, these bounds are independent of $d$, which explains the efficient PAC and online learnability of $\mathcal{H}_{\gamma}^{d}$ in arbitrarily high-dimensional spaces.
-Recent advances in learning theory, sparked by the influential works on the differential privacy of PAC learning [ALMM19, BLM20, ABL+22], have established that for total concept classes, finite Littlestone dimension not only characterizes online learnability but is also equivalent to list replicability, global stability, differential privacy (DP) learnability, and shared-randomness replicability.
-We will discuss these notions in depth shortly.
-This naturally leads to the question of whether these equivalences extend to partial concept classes.
-The primary goal of this paper is to provide a negative answer to this question by proving that $\mathcal{H}_{\gamma}^{\infty}$ is not list replicable, even though it has a bounded Littlestone dimension by Eq.
-2.
-Before delving deeper into these results, we formally define some key notions.
-Consider a partial concept class {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$.
-A distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ is realizable by a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ if, for every $n$, a random sample $\bm{S}=((\bm{x}_{i},\bm{y}_{i}))_{i=1}^{n}\sim\mathcal{D}^{n}$ is almost surely realizable by some $h\in\mathcal{H}$, meaning that $h(\bm{x}_{i})=\bm{y}_{i}$ for all $i=1,\ldots,n$.
-Note that $\mathcal{D}$ is a distribution over $\mathcal{X}\times\left\{\pm 1\right\}$, so none of the labels $\bm{y}_{i}$ take the value $\star$.
-Here, and throughout the paper, we use boldface letters to denote random variables and use the notation $(\bm{x},\bm{y})\sim\mathcal{D}$ to express that $(\bm{x},\bm{y})$ is a random variable distributed according to $\mathcal{D}$.
-The population loss of a partial concept $h\in\left\{\pm 1,\star\right\}^{\mathcal{X}}$ with respect to a distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ is
-$\mathcal{L}_{\mathcal{D}}(h)\coloneqq\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}[h(\bm{x})\neq\bm{y}].$
-Throughout this work, a learning rule refers to a (randomized) function $\bm{\mathcal{A}}$ that maps any sample $S\in\bigcup_{n=0}^{\infty}(\mathcal{X}\times\left\{\pm 1\right\})^{n}$ to a hypothesis $\bm{\mathcal{A}}(S)\in\left\{\pm 1\right\}^{\mathcal{X}}$.
-Since our primary focus is sample complexity rather than computational efficiency, we impose no computability constraints on $\bm{\mathcal{A}}$.
-Replicability.
-Replicability refers to the ability of an algorithm or study to produce consistent results when repeated under similar conditions and with similar data.
-It is a key principle of the scientific method that ensures research findings are reliable and not due to chance.
-In recent years, a rapidly growing body of research has emerged that introduced various rigorous formulations of replicability for learning algorithms, particularly in the context of the PAC learning framework [BLM20, MM22, CMY23, BGH+23, KVYZ23, EKK+23, EKM+23, MSS23, EHKS23, KKL+24, KKMV23].
-We introduce only the notions that are central to our main results and refer the reader to prior works, such as [MSS23, KKMV23, MM22], for a discussion of other related notions.
-Throughout, we use replicability as an umbrella term referring to the various formulations of this general concept.
-The concept of replicability in PAC learning first emerged in [BLM20, ABL+22] in the study of differential privacy of PAC learning algorithms.
-These works introduced a notion of replicability known as global stability to derive privacy guarantees from online learnability.
-A learning rule $\bm{\mathcal{A}}$ for a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is $(\epsilon,\rho)$-globally stable if for every realizable distribution $\mathcal{D}$, there is a hypothesis $h\in\left\{\pm 1\right\}^{\mathcal{X}}$ with population loss $\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon$ satisfying
-$\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})=h]\geq\rho,\ \ \text{where }n=n(\epsilon).$
-In other words, when run with samples from $\mathcal{D}$, there is a non-negligible chance $\rho>0$ that the learner will output the same hypothesis $h$.
-We define $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})$ to be the supremum of $\rho$ such that there is a $(\epsilon,\rho)$-globally stable learner for $\mathcal{H}$.
-The global stability parameter of $\mathcal{H}$ is then defined as
-$\rho^{\mathrm{gs}}(\mathcal{H})\coloneqq\inf_{\epsilon>0}\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H}).$
-The definition of global stability might initially seem weak, as a globally stable learner is not necessarily a PAC learner.
-In particular, since $\rho$ can be a small constant, there may be a probability as great as $1-\rho$ that the learning rule outputs a hypothesis with large population loss.
-However, [CMY23] showed that for total classes {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$, global stability is equivalent to a stronger notion called list replicability.
-Definition 1.1 (List replicability).
-A learning rule $\bm{\mathcal{A}}$ with sample complexity $n(\epsilon,\delta)$ is said to be an $(\epsilon,L)$-list replicable learner for {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$ if the following holds.
-For every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$, there is a list of total hypotheses $h_{1},\ldots,h_{L}\in\left\{\pm 1\right\}^{\mathcal{X}}$ such that
-  • The population loss of every $h_{i}$ with respect to $\mathcal{D}$ is at most $\epsilon$:
-    $\mathcal{L}_{\mathcal{D}}(h_{i})\leq\epsilon\text{ for every }i=1,\ldots,L.$
-  • The probability that $\bm{\mathcal{A}}$ outputs a hypothesis outside the list is at most $\delta$:
-    $\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})\not\in\left\{h_{1},\ldots,h_{L}\right\}]\leq\delta\ \text{ where }n=n(\epsilon,\delta).$
-Given an accuracy parameter $\epsilon\in(0,1)$, the $\epsilon$-list replicability number of $\mathcal{H}$, denoted by $\operatorname{LR}_{\epsilon}(\mathcal{H})$, is the smallest integer $L$, such that there exists an $(\epsilon,L)$-list replicable learner for $\mathcal{H}$.
-We define $\operatorname{LR}_{\epsilon}(\mathcal{H})=\infty$ if no such integer exists.
-The list replicability number of $\mathcal{H}$ is $\operatorname{LR}(\mathcal{H})\coloneqq\sup_{\epsilon>0}\operatorname{LR}_{\epsilon}(\mathcal{H}).$ We say $\mathcal{H}$ is list replicable if $\operatorname{LR}(\mathcal{H})<\infty$.
-Definition 1.1 provides a strong notion of replicability as the learner’s output is typically chosen from a small list $\left\{h_{1},\ldots,h_{L}\right\}$, and all these hypotheses have small population loss.
-In [CMY23], Chase, Moran and Yehudayoff proved that for every total class {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$, list replicability is equivalent to global stability.
-It is easy to check that their proof applies to partial concept classes, resulting in the following relationship between the list replicability number and global stability parameter.
-Theorem 1.2 ([CMY23]).
-For every partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, we have
-$\rho^{\mathrm{gs}}(\mathcal{H})=\frac{1}{\operatorname{LR}(\mathcal{H})}.$
-For completeness, we provide a proof in Appendix A.
-As shown in [BLM20], the global stability parameter of any nontrivial concept class is at most $1/2$.
-This fundamental limitation implies that no nontrivial class can achieve true replicability in the strongest sense, where the learning algorithm outputs the same hypothesis with high probability.
-To remedy this, Impagliazzo, Lei, Pitassi, and Sorrell [ILPS22] introduced a weaker notion of replicability where different executions of the algorithm can use the same random seed.
-Formally, let $\mathcal{A}(S,r)$ be a randomized learning rule, where $r$ denotes the random seed.
-Definition 1.3 (Shared-randomness replicability).
-A concept class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$ is shared-randomness replicable if there exists a learning rule $\mathcal{A}$ and a sample complexity function $n(\epsilon,\delta)$ such that, for every $\epsilon,\delta>0$ and every realizable distribution $\mathcal{D}$, the following conditions hold:
-  • Small population loss: $\Pr_{\bm{S}\sim\mathcal{D}^{n},\bm{r}}[\mathcal{L}_{\mathcal{D}}\left(\mathcal{A}(\bm{S},\bm{r})\right)>\epsilon]\leq\delta.$
-  • Replicability with shared randomness: $\Pr_{\bm{S,S^{\prime}}\sim\mathcal{D}^{n},\bm{r}}[\mathcal{A}(\bm{S},\bm{r})=\mathcal{A}(\bm{S^{\prime}},\bm{r})]\geq 1-\delta.$
-The following theorem, which brings together several celebrated recent results in learning theory, establishes that for total concept classes, all these notions of replicability, along with approximate differential privacy, are characterized by the finiteness of the Littlestone dimension.
-Theorem 1.4 ([ALMM19, BLM20, ABL+22, CMY23, ILPS22]).
-Let {±$\mathcal{H}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$ be a total concept class. The following statements are equivalent.
-  • $\operatorname{Ldim}(\mathcal{H})<\infty$.
-  • $\mathcal{H}$ is globally stable and equivalently list replicable.
-  • $\mathcal{H}$ is shared-randomness replicable.
-  • $\mathcal{H}$ is approximately differentially private (DP)-learnable (See Definition 2.9).
-1.1 Main contributions
-Replicability of large-margin half-spaces.
-The differentially private and replicable learning of large-margin half-spaces have been studied extensively [BDMN05, LNUZ20, BMNS19, KMST20, BCS20, BMS22a, BMS22b, ILPS22, KKL+24].
-In [LNUZ20], Nguyen, Ullman, and Zakynthinou gave a DP-learner for $\mathcal{H}^{d}_{\gamma}$ with a sample complexity that does not depend on the dimension $d$.
-In fact, they proved a stronger result that $\mathcal{H}^{d}_{\gamma}$ is pure DP-learnable (Definition 2.10).
-Combining the dimension-free DP-learnability of $\mathcal{H}^{d}_{\gamma}$ with the “DP-learnable to Shared-randomness Replicability” reduction of [BGH+23] shows that $\mathcal{H}^{d}_{\gamma}$ is shared-randomness replicable with dimension-free sample complexity.
-More recently, Kalavasis, Karbasi, Larsen, Velegkas, and Zhou [KKL+24] employed a more sophisticated rounding scheme to directly obtain a shared-randomness replicability of this problem, which further improves the sample complexity.
-In the following theorem, we summarize the consequences of these lines of work that are relevant to this paper.
-Theorem 1.5 ([MP43, Ros58, LNUZ20, KKL+24]).
-For every $\gamma>0$ and $d\in\mathbb{N}\cup\left\{\infty\right\}$, the class $\mathcal{H}^{d}_{\gamma}$ satisfies the following.
-  • $\operatorname{Ldim}(\mathcal{H}^{d}_{\gamma})<\gamma^{-2}$.
-  • The class $\mathcal{H}^{d}_{\gamma}$ is (pure) DP-learnable with dimension-independent sample complexity.
-  • The class $\mathcal{H}^{d}_{\gamma}$ is shared-randomness replicable with dimension-independent sample complexity.
-Given the positive results in Theorem 1.5 and the equivalences in Theorem 1.4, one might expect that, as in the case of total concepts, the list replicability number of $\mathcal{H}^{d}_{\gamma}$ is independent of $d$.
-Our main theorem establishes that this is not the case—the list replicability number of $\mathcal{H}^{d}_{\gamma}$ grows as $d$ increases.
+Gap Hamming problem.
+The discrete analogue of large-margin half-spaces is the well-studied Gap Hamming Distance (GHD) problem, a central problem in communication complexity.
+For $n\in\mathbb{N}$ and $\gamma\in(0,1)$, the $n$-bit $\operatorname{GHD}_{\gamma}$ problem, denoted $\operatorname{GHD}_{\gamma}^{n}$, is the partial function on inputs $x,y\in\{\pm 1\}^{n}$ defined by
+$\operatorname{GHD}_{\gamma}^{n}(x,y)\coloneqq\begin{cases}\operatorname{sgn}({\left\langle x,y\right\rangle})&\text{if $|{\left\langle x,y\right\rangle}|>\gamma n$}\\ \star&\text{otherwise}\end{cases}.$
+As a communication problem, Alice receives $x$, Bob receives $y$, and under the promise $|{\left\langle x,y\right\rangle}|\geq\gamma n$, they must compute $\operatorname{GHD}_{\gamma}^{n}(x,y)$ with minimal communication.
+For fixed $\gamma$, the public-coin randomized communication complexity of $\operatorname{GHD}_{\gamma}^{n}$ is $O_{\gamma}(1)$: using shared randomness, the players sample a subset $S\subseteq[n]$ of size $O_{\gamma}(1)$ and estimate $\left\langle x,y\right\rangle$} by $\tfrac{n}{|S|}\sum_{i\in S}x_{i}y_{i}$, which requires only $2|S|$ communicated bits.
+2 Main theorem
+Our main theorem determines the list replicability number of $\mathcal{H}_{\gamma}^{d}$ up to a factor of two, showing in particular that it grows unboundedly with the dimension $d$.
 Theorem 2.1 (Main theorem).
-For any fixed margin $\gamma\in(0,1)$, finite dimension $d>1$, and accuracy parameter $\epsilon\in(0,1/2)$,
+For any fixed margin $\gamma\in(0,1)$, dimension $d>1$, and accuracy parameter $\epsilon\in(0,1/2)$,
 $\frac{d}{2}+1\leq\operatorname{LR}_{\epsilon}(\mathcal{H}_{\gamma}^{d})\leq d.$
-In particular, $\mathcal{H}_{\gamma}^{\infty}$ is not list replicable (equiv. globally stable).
-Theorem 1.6 reveals a surprising distinction in the partial setting: list replicability does not follow from bounded Littlestone dimension, replicability, DP-learnability, or even pure DP-learnability.
-Specifically, for every $\gamma\in(0,1)$, the partial concept class $\mathcal{H}^{\infty}_{\gamma}$ is shared-randomness replicable, pure DP-learnable, and has a finite Littlestone dimension, but it is not list replicable.
+Hence, $\tfrac{d}{2}+1\leq\operatorname{LR}(\mathcal{H}_{\gamma}^{d})\leq d$.
 The lower bound in Theorem 2.1 relies on a topological argument involving covers of the sphere by antipodal-free open sets.
 In particular, we apply the local Borsuk-Ulam theorem of [CCMY24], which states that in such a cover, there is a point that belongs to at least $\frac{d}{2}+1$ sets.
-Alternatively, one could use Ky Fan’s classical theorem [Fan52], but this would yield the slightly weaker bound of $\frac{d}{2}$.
+Alternatively, one could use Ky Fan’s classical theorem [Fan52], but this would yield the slightly weaker lower bound of $\frac{d}{2}$.
 For the upper bound, we construct a learning rule that uses the generalization properties of hard-SVM combined with a list-replicable rounding scheme using a fine net in general position.
-Disambiguations of Large-margin Half-spaces:
-One of the central questions studied in [AHHM22] is whether the learnability of a partial concept class can always be inferred from the learnability of a suitable extension of it to a total concept class.
-Formally, a disambiguation of a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is a total concept class ¯{±$\overline{\mathcal{H}}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$ such that for every $h\in\mathcal{H}$ and every finite $S\subseteq h^{-1}(\left\{\pm 1\right\})$, there exists an $\bar{h}\in\overline{\mathcal{H}}$ that is consistent with $h$ on $S$.
-Intuitively, one may think of disambiguation as replacing each $\star$ with $-1$ or $+1$, although this intuition is not completely rigorous in the infinite case.
-Note that any distribution realizable by a partial concept class $\mathcal{H}$ is also realizable by any disambiguation $\overline{\mathcal{H}}$ of $\mathcal{H}$.
-Consequently, if a disambiguation $\overline{\mathcal{H}}$ is efficiently learnable, the same holds true for $\mathcal{H}$.
-Motivated by the fact that $\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})\leq\gamma^{-2}$, [AHHM22] posed the following question.
-Question 1.7 ([AHHM22, Question 4]).
-Does there exist a disambiguation of $\mathcal{H}_{\gamma}^{d}$ by a total concept class whose VC/Littlestone dimension is bounded by a function of $\gamma$?
-In the following theorem, we provide a negative answer to the Littlestone dimension part of 1.7.
+2.1 Applications
+Separation.
+In addition to list replicability, differentially private learnability and shared-randomness replicability [^2] In prior works, shared-randomness replicability is referred to simply as replicability.
+Since we work with multiple replicability notions, we adopt this terminology to emphasize that different executions of the algorithm use the same random seed.
+are two other well-studied notions of stability in learning theory (see [MSS23] for an overview).
+For completeness, formal definitions of these concepts appear in the appendix, though we do not rely on them directly in this paper.
+Recent advances in learning theory [ALMM19, BLM20, ABL+22, CMY23, ILPS22], sparked by the influential works on differential privacy in PAC learning, have established that for total concept classes, all of these notions coincide and are characterized by the finiteness of the Littlestone dimension.
+Specifically, for every total concept class $\mathcal{C}\subseteq\{\pm 1\}^{\mathcal{X}}$, the following are equivalent:
+  • $\operatorname{Ldim}(\mathcal{C})<\infty$;
+  • $\mathcal{C}$ is list replicable;
+  • $\mathcal{C}$ is shared-randomness replicable;
+  • $\mathcal{C}$ is approximately differentially private (DP)-learnable.
+This naturally raises the question of whether these equivalences extend to partial concept classes.
+The case of large-margin half-spaces has been studied extensively [BDMN05, LNUZ20, BMNS19, KMST20, BCS20, BMS22a, BMS22b, ILPS22, KKL+24], and the following are known:
+  • $\operatorname{Ldim}(\mathcal{H}^{d}_{\gamma})<\gamma^{-2}$;
+  • $\mathcal{H}^{d}_{\gamma}$ is (pure) DP-learnable with dimension-independent sample complexity;
+  • $\mathcal{H}^{d}_{\gamma}$ is shared-randomness replicable with dimension-independent sample complexity.
+Nevertheless, Theorem 2.1 shows that despite these strong positive results, the list replicability number of $\mathcal{H}^{d}_{\gamma}$ necessarily grows with $d$.
+Consequently, we obtain a sharp separation from the total setting: for partial classes, list replicability does not follow from bounded Littlestone dimension, replicability, DP-learnability, or even pure DP-learnability.
+Corollary 2.2 (Separation).
+There exists a partial concept class $\mathcal{C}$ that is (pure) DP-learnable, shared-randomness replicable, and satisfies $\operatorname{Ldim}(\mathcal{C})<\infty$, yet it is not list replicable.
+Proof.
+Fix $\gamma\in(0,1)$, and define the class $\mathcal{H}^{\infty}_{\gamma}$ as follows.
+Each hypothesis in $\mathcal{H}^{\infty}_{\gamma}$ is specified by a unit vector $w$ of arbitrary finite dimension, i.e., $w\in\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}$.
+For $x\in\bigcup_{d\in\mathbb{N}}\mathbb{S}^{d-1}$, define
+$c_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}({\left\langle w,x\right\rangle})&\text{if $\dim(x)=\dim(w)$ and $|{\left\langle w,x\right\rangle}|\geq\gamma$,}\\ \star&\text{otherwise}\\ \end{cases}.$
+By the aforementioned results of [MP43, Ros58, LNUZ20, KKL+24], the class $\mathcal{H}^{\infty}_{\gamma}$ is pure DP-learnable, shared-randomness replicable, and satisfies $\operatorname{Ldim}(\mathcal{H}^{\infty}_{\gamma})<\gamma^{-2}$.
+However, by Theorem 2.1, we have $\operatorname{LR}(\mathcal{H}^{\infty}_{\gamma})=\infty$.
+This establishes the claim.
+∎
+Disambiguations of large-margin half-spaces.
+A disambiguation of a partial concept class $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is a total concept class $\overline{\mathcal{C}}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$ such that for every $c\in\mathcal{C}$ and every finite $S\subseteq c^{-1}(\left\{\pm 1\right\})$, there exists an $\bar{c}\in\overline{\mathcal{C}}$ that is consistent with $c$ on $S$.
+Intuitively, this corresponds to resolving each $\star$ with 1-1 or +1+1, although this intuition is not completely rigorous in the infinite case.
+Disambiguation cannot decrease the list-replicability number.
+At the same time, it converts a partial class into a total class, where the highly nontrivial results of [BLM20, ABL+22, GGKM21] show that list replicability is bounded in terms of the Littlestone dimension.
+This principle underlies our results on disambiguations of the large-margin half-space problem and its discrete analogue $\operatorname{GHD}_{\gamma}$.
+The following theorem resolves an open question of [AHHM21, Question 4] who asked whether every disambiguation of $\mathcal{H}_{\gamma}^{d}$ satisfies $\operatorname{Ldim}=\omega(1)$.
 Theorem 2.3.
@@
 Proof.
-This follows by combining the lower bound of Theorem 1.6 and a result of [GGKM21] showing that for every total concept class $\mathcal{H}$, $\operatorname{LR}_{\epsilon}(\mathcal{H})\leq 2^{O_{\epsilon}\left(\operatorname{Ldim}(\mathcal{H})^{2}\right)}$.
+As was noted in [CMY23], it is implicitly  [^3] See [GGKM21, Lemma 5.5] and the definitions of $k_{t}$ and $k^{\prime}$, where $k^{\prime}$ depends on $n_{0}$, $d_{L}$, $\alpha_{\Delta}$, $C_{0}$, and $\eta^{2}$ as specified in [GGKM21, Algorithm 2].
+proved in [GGKM21] that for every class $\overline{\mathcal{H}}$,
+$\operatorname{LR}_{\epsilon}(\overline{\mathcal{H}})\leq 2^{O_{\epsilon}\left(\operatorname{Ldim}(\overline{\mathcal{H}})^{2}\right)}.$
+Recall that by Theorem 2.1, for every $\epsilon\in(0,1/2)$, $\operatorname{LR}_{\epsilon}(\overline{\mathcal{H}})\geq\frac{d}{2}+1$.
+Combining the two inequalities for a fixed $\epsilon\in(0,1/2)$ concludes the claim.
 ∎
-On the other hand, there is a disambiguation $\overline{\mathcal{H}}$ of $\mathcal{H}_{\gamma}^{d}$ with Littlestone dimension $\operatorname{Ldim}(\overline{\mathcal{H}})=O(d)$ since [HHM23] shows that there is a disambiguation of $\mathcal{H}_{\gamma}^{d}$ into a finite total concept class $\overline{\mathcal{H}}$ with $2^{O(d)}$ concepts.
-The upper bound on the Littlestone dimension follows from the fact that every finite class $\mathcal{H}$ satisfies $\operatorname{VCdim}(\mathcal{H})\leq\operatorname{Ldim}(\mathcal{H})\leq\log|\mathcal{H}|$.
-Theorem 1.8 has two important implications.
-Firstly, it shows $\mathcal{H}^{\infty}_{\gamma}$ is a partial class with finite Littlestone dimension, but any disambiguation of it has an infinite Littlestone dimension.
-The existence of such partial classes was posed as a question in [AHHM22] and answered in [CHHH23] through a complex construction based on Göös’ breakthrough [Göö15] in communication complexity.
-Theorem 1.8 offers a more natural example of a concept class with this property.
-Secondly, the construction of  [CHHH23] has Littlestone dimension $2$, and in [CHHH23, Question 4.1], they asked whether there exists a partial concept class with Littlestone dimension $1$ such that all its disambiguations have infinite Littlestone dimension.
-The following corollary of Theorem 1.8 resolves their question.
+[AHHM21] used a sophisticated construction, building on Göös’ breakthrough refutation of the Alon–Saks–Seymour conjecture [Göö15], to exhibit partial concept classes with $\operatorname{VCdim}=2$ whose disambiguations satisfy $\operatorname{Ldim}=\omega(1)$.
+Subsequently, [CHHH23] employed a similar approach to construct partial classes with $\operatorname{Ldim}=2$ whose disambiguations again satisfy $\operatorname{Ldim}=\omega(1)$.
+Theorem 2.3 provides a much more natural example of a class exhibiting this phenomenon.
+Moreover, [CHHH23, Question 4.1] asked whether such a separation can already occur for partial classes of Littlestone dimension $1$.
+The following corollary of Theorem 2.3 answers this question in the affirmative.
 Corollary 2.4.
-Let $\gamma\in(1/\sqrt{2},1)$. Then $\operatorname{Ldim}(\mathcal{H}_{\gamma}^{\infty})=1$, while any disambiguation of $\mathcal{H}_{\gamma}^{\infty}$ has infinite Littlestone dimension.
+Let $\gamma\in(1/\sqrt{2},1)$. Then $\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})=1$, while any disambiguation of $\mathcal{H}_{\gamma}^{d}$ has Littlestone dimension $\Omega(\sqrt{\log d})$.
 Proof.
-We have $21\leq\operatorname{Ldim}(\mathcal{H}_{\gamma}^{\infty})\leq 1/\gamma^{2}<$ 2. An application of Theorem 1.8 completes the proof. ∎
-List replicability of finite classes.
-In [CMY23, Theorems 5 and 13], Chase, Moran, and Yehudayoff proved that the list replicability number of any finite set of homogeneous half-spaces and points that do not lie on the half-spaces in $\mathbb{R}^{2}$ is at most $2$.
-They further asked whether a similar finite bound holds in higher dimensions.
-As we discuss below, we give an affirmative answer to their question in Theorem 1.11.
-Definition 1.10.
-Define the finitary list replicability number of a concept class {±$\mathcal{H}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$, denoted $\widetilde{\operatorname{LR}}(\mathcal{H})$, as
-$\widetilde{\operatorname{LR}}(\mathcal{H})\coloneqq\sup_{\text{finite }S\subseteq\mathcal{X}}\operatorname{LR}(\mathcal{H}|_{S}).$
-Given $d\in\mathbb{N}$, let $\mathcal{H}^{d}$ be the partial concept class defined by points in $\mathbb{S}^{d-1}$ and partial concepts $h_{w}:\mathbb{S}^{d-1}\to\left\{\pm 1,\star\right\}$, where $w\in\mathbb{S}^{d-1}$, given by
-$h_{w}(x)\coloneqq\begin{cases}\operatorname{sgn}(\langle w,x\rangle)&\text{if }\langle w,x\rangle\neq 0\\ \star&\text{otherwise}\end{cases}.$
-Note that for every $\gamma>0$, $\mathcal{H}^{d}_{\gamma}$ is a subclass of $\mathcal{H}^{d}$.
-Since $\mathcal{H}^{2}$ has infinite threshold dimension [AHHM22, Definition 28], by the celebrated result of [ABL+22], it is not globally stable and therefore $\operatorname{LR}(\mathcal{H}^{2})=\infty$.
-Consequently, the aforementioned result of [CMY23] implies the following interesting gap:
+We have $1\leq\operatorname{Ldim}(\mathcal{H}_{\gamma}^{d})\leq 1/\gamma^{2}<2$. An application of Theorem 2.3 completes the proof. ∎
+Disambiguations of gap Hamming distance.
+In complexity theory, separations between complexity measures are often easier to demonstrate for partial functions, and disambiguations of important partial functions are studied as a way to extend such results to the total setting.
+For fixed $\gamma\in(0,1)$ the partial $\operatorname{GHD}_{\gamma}$ is known to separate constant cost randomized communication complexity from several other important complexity measures [HHH23, CLV19, HHM23, Son14].
+Motivated by this, researchers have asked whether $\operatorname{GHD}_{\gamma}$ admits a disambiguation with constant cost randomized communication complexity [FGHH25].
+Our next theorem gives a negative answer to this question.
+Theorem 2.5.
+Let $\gamma\in(0,1)$ be a margin parameter. Every family of disambiguations $\{M_{n}\}_{n=1}^{\infty}$ of the Gap Hamming Distance matrices $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$ satisfies
+$\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n}),$ (3)
+and has public-coin randomized communication complexity $\Omega(\log\log n)$.
+To lower bound the Littlestone dimension, we use an embedding due to [HHM23] that allows us to disambiguate $\mathcal{H}_{\gamma}^{d}$ using a disambiguation of the Gap Hamming Distance problem in dimension $O(d)$.
+The key here is that the embedding will maintain the lower bound on Littlestone dimension.
+Equation 3 then follows from Theorem 2.3.
+The bound on the communication complexity then follows as a corollary, using the known relationship between Littlestone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity.
+See Section 4 for the proof.
+Pseudo-determinism vs randomness in communication.
+A pseudo-deterministic algorithm is a randomized algorithm that, when executed multiple times on the same input, produces the same output with high probability.
+This notion was introduced by Gat and Goldwasser [GG11] and has since been extensively investigated across a variety of computational models, including learning algorithms, communication protocols, decision tree algorithms, sequential and parallel algorithms, average-case and approximation algorithms, interactive proofs, low-space algorithms, and streaming algorithms (see [GGMW20, GIPS21] and the references therein).
+A central question in this line of research is to understand to what extent pseudo-determinism can be separated from general randomized computation.
+In communication complexity, a search problem is specified by a relation $\mathcal{R}\subseteq\mathcal{X}\times\mathcal{Y}\times\mathcal{Z}$, where Alice and Bob receive $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ respectively, and must output $z\in\mathcal{Z}$ such that $(x,y,z)\in\mathcal{R}$ while minimizing communication.
+In the public-coin randomized model, the players have access to a shared random string $\bm{r}$, and a protocol $\pi$ must satisfy
+$\Pr_{\bm{r}}[(x,y,\pi(\bm{r},x,y))\in\mathcal{R}]\geq\frac{2}{3}\ \ \forall(x,y)\in\mathcal{X}\times\mathcal{Y}.$
+Such a protocol is called pseudo-deterministic if there exists a function $f:\mathcal{X}\times\mathcal{Y}\to\mathcal{Z}$ (with $(x,y,f(x,y))\in\mathcal{R}$ for all $(x,y)$) such that
+$\Pr_{\bm{r}}[\pi(\bm{r},x,y)=f(x,y)]\geq\frac{2}{3}\ \ \forall(x,y)\in\mathcal{X}\times\mathcal{Y}.$
+A well-known candidate for separating pseudo-determinism from randomized communication is the approximate Hamming distance problem
+$(x,y,t)\in\operatorname{AHD}_{n}\ \Longleftrightarrow\ |d_{H}(x,y)-t|<\tfrac{n}{3},$
+where $d_{H}(x,y)$ is the Hamming distance between $x,y\in\{\pm 1\}^{n}$.
+In the public-coin randomized model, Alice and Bob can solve $\operatorname{AHD}_{n}$ with only $O(1)$ bits of communication by sampling $O(1)$ coordinates uniformly at random and exchanging the corresponding entries to estimate $d_{H}(x,y)$.
+By contrast, it is widely believed that the pseudo-deterministic communication complexity of $\operatorname{AHD}_{n}$ is large.
+The following theorem establishes the first super-constant lower bound on this problem, thereby yielding the first $O(1)$ versus $\omega(1)$ separation between randomized and pseudo-deterministic communication complexities.
+Theorem 2.6.
+For any $\epsilon<\frac{1}{2}$, the pseudo-deterministic communication complexity of $\operatorname{AHD}_{n}$ is $\Omega(\log\log(n))$.
+Proof.
+Suppose there is a pseudo-deterministic protocol for $\operatorname{AHD}_{n}$ of cost $k$, with corresponding function $f:\{\pm 1\}^{n}\times\{\pm 1\}^{n}\to\{0,1,\ldots,n\}$. Define
+$F(x,y)\coloneqq\begin{cases}1&f(x,y)\geq\frac{n}{2}\\ -1&f(x,y)<\frac{n}{2}\end{cases},$
+and note that the public-coin randomized communication complexity of $F$ is at most $k$.
+Since $f(x,y)$ is always within $n/3$ of the true Hamming distance $d_{H}(x,y)$, the function $F$ disambiguates the Gap Hamming Distance problem $\operatorname{GHD}_{0.1}^{n}$, and by Theorem 2.3, its randomized communication complexity is $\Omega(\log\log n)$.
+∎
+List replicability of finite hyperplane arrangements.
+Define the finitary list replicability number of a concept class $\mathcal{C}\subseteq\{\pm 1,\star\}^{\mathcal{X}}$ as
+$\widetilde{\operatorname{LR}}(\mathcal{C})\coloneqq\sup_{\text{finite }S\subseteq\mathcal{X}}\operatorname{LR}(\mathcal{C}|_{S}).$
+As an example, consider the class $\mathcal{H}^{2}$ of homogeneous halfspaces in $\mathbb{R}^{2}$.
+While $\operatorname{LR}(\mathcal{H}^{2})=\infty$, Chase, Moran, and Yehudayoff [CMY23, Theorems 5 and 13] proved that for any finite set of points $S\subseteq\mathbb{S}^{1}$, we have $\operatorname{LR}(\mathcal{H}^{2}_{0}|_{S})\leq 2$.
+This establishes the following striking gap:
 $\operatorname{LR}(\mathcal{H}^{2})=\infty\qquad\text{while}\qquad\widetilde{\operatorname{LR}}(\mathcal{H}^{2})=2.$
-The authors of [CMY23] ask whether a similar bound holds for $\widetilde{\operatorname{LR}}(\mathcal{H}^{3})$.
-In the following theorem, we answer their question by extending their result to all dimensions.
+They further asked whether a similar bound holds for $\widetilde{\operatorname{LR}}(\mathcal{H}^{3})$ and, more generally, in higher dimensions. The next theorem resolves their question affirmatively.
 Theorem 2.7.
@@
 The upper bound is an easy consequence of the upper bound of our main theorem (Theorem 2.1).
-Indeed, any finite set of points $S\subset\mathbb{S}^{d-1}$ and hypotheses $H\subset\mathcal{H}^{d}|_{S}$ defined by unit vectors $W\subset\mathbb{S}^{d-1}$ is a sub-concept class of $\mathcal{H}^{d}_{\gamma}$, where $\gamma=\min_{x\in S,w\in W}|\langle w,x\rangle|$.
+Indeed, any finite set of points $S\subset\mathbb{S}^{d-1}$ and hypotheses $H\subset\mathcal{H}^{d}|_{S}$ defined by unit vectors $W\subset\mathbb{S}^{d-1}$ is a sub-concept class of $\mathcal{H}^{d}_{\gamma}$, where $\gamma\coloneqq\min_{x\in S,w\in W}|{\left\langle w,x\right\rangle}|$.
 The claim now follows from our upper bound from Theorem 2.1.
-For the lower bound, we use a result of Chase, Moran, and Yehudayoff [CMY23, Theorem 3] stating that for every concept class $\mathcal{H}$,
-$\operatorname{LR}(\mathcal{H})\geq\operatorname{VCdim}(\mathcal{H}).$
-The result now follows as $\operatorname{VCdim}(\mathcal{H}^{d})=d$, and therefore, $\mathcal{H}^{d}$ has a finite subclass of VC dimension $d$. ∎
-The sign-rank of a sign matrix $A$, denoted by ${\mathbf{rk}}_{\pm}(A)$, is the smallest rank of a real matrix $B_{\mathcal{X}\times\mathcal{Y}}$ such that the entries of $B$ are nonzero and have the same signs as their corresponding entries in $A$.
-The definition of sign-rank naturally extends to partial matrices, where for invalid entries of $A$, the corresponding entry in $B$ could be any real number.
+For the lower bound, we use a result of Chase, Moran, and Yehudayoff [CMY23, Theorem 3] stating that for every concept class $\mathcal{C}$,
+$\operatorname{LR}(\mathcal{C})\geq\operatorname{VCdim}(\mathcal{C}).$
+The result follows as $\operatorname{VCdim}(\mathcal{H}^{d})=d$, and hence, $\mathcal{H}^{d}$ has a finite subclass of VC dimension $d$. ∎
+Theorem 2.7 reveals a connection between list replicability and one of the most fundamental parameters in learning theory called sign-rank.
 Geometrically, sign-rank is the smallest dimension in which the matrix is realized as points and homogeneous half-spaces.
-We will state the definition of sign-rank in this terminology for partial classes.
 Definition 2.8 (Sign-rank).
-The sign-rank of a partial class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, denoted by ${\mathbf{rk}}_{\pm}(\mathcal{H})$, is the smallest $d$ such that there exist vectors $u_{h},v_{x}\in\mathbb{R}^{d}$ for all pairs $h\in\mathcal{H},x\in\mathcal{X}$ such that $h(x)=\operatorname{sgn}({\left\langle u_{h},v_{x}\right\rangle})$ whenever $h(x)\neq\star$.
-Theorem 1.11 and the VC lower bound of [CMY23, Theorem 3] immediately imply the following general bounds on $\widetilde{\operatorname{LR}}(\mathcal{H})$.
+The sign-rank of a partial class $\mathcal{C}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, denoted by $\operatorname{sign-rank}(\mathcal{C})$, is the smallest $d$ such that there exist vectors $u_{c},v_{x}\in\mathbb{R}^{d}$ for all pairs $c\in\mathcal{C},x\in\mathcal{X}$ such that $c(x)=\operatorname{sgn}({\left\langle u_{c},v_{x}\right\rangle})$ whenever $c(x)\neq\star$.
+Combining Theorem 2.7 with the VC-dimension lower bound of [CMY23, Theorem 3] yields the following general bounds on the finitary list replicability number.
 Corollary 2.9.
-For every partial class {±,}$\mathcal{H}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, we have
-$\operatorname{VCdim}(\mathcal{H})\leq\widetilde{\operatorname{LR}}(\mathcal{H})\leq\operatorname{rank_{\pm}}(\mathcal{H}).$
-Gap Hamming Distance.
-The discrete analogue of large-margin half-spaces is the well-studied Gap Hamming Distance (GHD) problem, a central problem in communication complexity.
-For $n\in\mathbb{N}$ and $\gamma\in(0,1)$, the $n$-bit $\operatorname{GHD}_{\gamma}$ problem, denoted $\operatorname{GHD}_{\gamma}^{n}$, is the partial function on inputs $x,y\in\{\pm 1\}^{n}$ defined by
-$\operatorname{GHD}_{\gamma}^{n}(x,y)\coloneqq\begin{cases}\operatorname{sgn}({\left\langle x,y\right\rangle})&\text{if $|{\left\langle x,y\right\rangle}|\geq\gamma n$}\\ \star&\text{otherwise}\end{cases}.$
-For fixed $\gamma\in(0,1)$, the public-coin randomized communication complexity of $\operatorname{GHD}_{\gamma}$ is at most a fixed constant depending only on $\gamma$.
-This places $\operatorname{GHD}_{\gamma}$ in the complexity class $\sf{BPP}_{0}$, which consists of communication problems with constant-cost public-coin randomized protocols.
-Moreover, $\operatorname{GHD}_{\gamma}$ exhibits certain properties that separate $\sf{BPP}_{0}$ from some complexity classes (e.g., $\sf{UPP}_{0}$) and measures [HHH23, CLV19, HHM23, FGHH25, Son14].
-A key caveat of these separations is that $\operatorname{GHD}_{\gamma}$ is a partial function, and whether it extends to any family of total functions with similar properties remained an open problem before this work.
-One suggested approach [FGHH25] to resolving these questions is to find a disambiguation of $\operatorname{GHD}_{\gamma}$ to a total function with constant public-coin randomized communication complexity.
-In Corollary 1.15 below, we prove this is impossible: every disambiguation of $\operatorname{GHD}_{\gamma}^{n}$ has public-coin randomized communication complexity at least $\Omega(\log\log n)$.
-First, we show that every such disambiguation has an unbounded Littlestone dimension.
-Theorem 1.14.
-Let $\gamma\in(0,1)$ be a margin parameter. Every family of disambiguations $\{M_{n}\}_{n=1}^{\infty}$ of the Gap Hamming Distance matrices $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$ satisfies
-$\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n}).$
-We obtain the following corollary by combining Theorem 1.14 with some known results.
-Corollary 1.15.
-Let $\gamma\in(0,1)$ be a margin parameter.
-Every family of disambiguation of the Gap Hamming Distance matrices $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$ has public-coin randomized communication complexity $\Omega(\log\log n)$.
-The bound follows by combining Theorem 1.14 with the known relationship between Littlestone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity.
-See Section 4 for the proof.
+For every partial class $\mathcal{C}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$, we have
+$\operatorname{VCdim}(\mathcal{C})\leq\widetilde{\operatorname{LR}}(\mathcal{C})\leq\operatorname{sign-rank}(\mathcal{C}).$
 2.2 Concluding remarks and open problems
-While Theorem 1.4 provides an elegant and comprehensive characterization of various notions of replicability and privacy for total concepts through the combinatorial framework of the Littlestone dimension, the landscape of replicability for partial classes, as demonstrated by the results of this paper, is more intricate and less understood.
+For total concept classes, list replicability, shared randomness replicability, and (approximate) DP-learnability are now known to coincide through the combinatorial framework of the Littlestone dimension.
+In contrast, the situation for partial classes, as illustrated by the results of this paper, is more intricate and less understood.
 The “DP-learnability to Shared-randomness replicability” reduction from [BGH+23] extends to the partial setting.
-Moreover, [FHM+24] recently showed that for partial classes, DP-learnability implies a finite Littlestone dimension.
-Theorem 1.16 ([BGH+23, FHM+24]).
-Let {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class.
-  • If $\mathcal{H}$ is DP-learnable, then $\mathcal{H}$ is shared-randomness replicable.
-  • If $\mathcal{H}$ is DP-learnable, then $\operatorname{Ldim}(\mathcal{H})<\infty$.
+Moreover, [FHM+24] recently showed that for partial classes, DP-learnability implies a finite Littlestone dimension, and [KKMV23, Lemma 8] shows that even in the partial setting, list replicability implies shared-randomness replicability.
+Theorem 2.10 ([BGH+23, FHM+24, KKMV23]).
+Let $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class.
+  • If $\mathcal{C}$ is DP-learnable, then $\mathcal{C}$ is shared-randomness replicable.
+  • If $\mathcal{C}$ is DP-learnable, then $\operatorname{Ldim}(\mathcal{C})<\infty$.
+  • If $\mathcal{C}$ is list replicable, then $\mathcal{C}$ is shared-randomness replicable.
 On the other hand, our main theorem shows that for partial concepts, list replicability does not follow from bounded Littlestone dimension, shared-randomness replicability, DP-learnability, or even pure DP-learnability.
-It is also straightforward to show that, even in the partial setting, list replicability implies shared-randomness replicability.
-Proposition 1.17 ([KKMV23, Lemma 8]).
-If {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is list replicable (equiv. globally stable), then $\mathcal{H}$ is shared-randomness replicable.
-To our knowledge, no further relationships among the four notions of DP-learnability, shared-randomness replicability, Littlestone dimension, and list replicability are known for partial functions.
-Open problems.
+To our knowledge, no further relationships among DP-learnability, shared-randomness replicability, Littlestone dimension, and list replicability are currently known for partial concept classes.
 Finally, we list some open problems for future research that naturally arise from our work.
@@
   2. 2.
-    Is there a combinatorial notion of dimension that characterizes list replicability?
+    Is there a simple combinatorial notion of dimension that characterizes list replicability?
   3. 3.
     How tight are the inequalities in Corollary 2.9, namely,
-    $\operatorname{VCdim}(\mathcal{H})\leq\widetilde{\operatorname{LR}}(\mathcal{H})\leq\operatorname{rank_{\pm}}(\mathcal{H})?$
+    $\operatorname{VCdim}(\mathcal{C})\leq\widetilde{\operatorname{LR}}(\mathcal{C})\leq\operatorname{sign-rank}(\mathcal{C})?$
+    Is it possible to upper-bound $\widetilde{\operatorname{LR}}(\mathcal{C})$ by a function of $\operatorname{VCdim}(\mathcal{C})$?
   4. 4.
-    Is the answer to 1.7 negative for VC dimension?
-    In an article posted on arXiv just a few days ago, Chornomaz, Moran, and Waknine [CMW25] explored this problem using a topological approach, but the question remains open.
-2 Preliminaries
-Much of the notation and definitions were already outlined in the introduction.
-In this section, we review a few key concepts from learning theory, specifically PAC learnability, the VC and Littlestone dimensions, and differential privacy.
-In some of our proofs, we will apply Hoeffding’s classic concentration inequality, which we state below for reference.
-Theorem 2.1 (Hoeffding’s inequality).
-Let $c\in\mathbb{R}$ and let $\bm{x}_{1},\ldots,\bm{x}_{n}$ be independent random variables with $\bm{x}_{i}\in[-c,c]$ and $\mathbb{E}[\bm{x}_{i}]=0$. Then, for any $t>0$,
-$\Pr\left[\left|\sum_{i=1}^{n}\bm{x}_{i}\right|\geq t\right]\leq 2e^{-\frac{t^{2}}{2c}}.$
-We define the population loss of a partial concept $h\in\left\{\pm 1,\star\right\}^{\mathcal{X}}$ with respect to a distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ as
-$\mathcal{L}_{\mathcal{D}}(h)\coloneqq\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}[h(\bm{x})\neq\bm{y}].$
-Under this definition, predictions with $h(x)=\star$ always count as mistakes. The population loss of a class $\mathcal{H}$ is given by the smallest population loss achievable within $\mathcal{H}$:
-$\mathcal{L}_{\mathcal{D}}(\mathcal{H})\coloneqq\inf_{h\in\mathcal{H}}\mathcal{L}_{\mathcal{D}}(h).$
-The following simple lemma from [AHHM22] establishes the connection between realizability and having zero population loss.
-Lemma 2.2 ([AHHM22]).
-Let {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class, and let $\mathcal{D}$ be a distribution on $\mathcal{X}\times\left\{\pm 1\right\}$.
-If $\mathcal{L}_{\mathcal{D}}(\mathcal{H})=0$, then $\mathcal{D}$ is realizable by $\mathcal{H}$.
-Conversely, if $\mathcal{D}$ is realizable and has finite or countable support, then $\mathcal{L}_{\mathcal{D}}(\mathcal{H})=0$.
-PAC learning of partial classes:
-Given a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, the goal of PAC (Probably Approximately Correct) learning is for the learning rule to produce, with high probability, a hypothesis whose population loss is close to the best achievable loss within $\mathcal{H}$.
-Definition 2.3 (PAC learning of partial concept classes).
-A class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is PAC learnable if there is a learning rule $\bm{\mathcal{A}}$ and a function $n(\epsilon,\delta)$ such that for every $\epsilon,\delta>0$ and every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm\right\}$, we have
-$\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\mathcal{L}_{\mathcal{D}}(\bm{\mathcal{A}}(\bm{S}))\leq\epsilon]\geq 1-\delta\ \text{ where }n=n(\epsilon,\delta).$
-2.1 VC and Littlestone dimensions
-We now extend the definition of the VC dimension to partial concept classes.
-Definition 2.4 (Shattered set).
-A finite set of points $C=\left\{x_{1},\ldots,x_{n}\right\}\subseteq\mathcal{X}$ is shattered by a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ if, for every pattern $y\in\left\{\pm 1\right\}^{n}$, there exists $h\in\mathcal{H}$ such that $h(x_{i})=y_{i}$ for all $i\in[n]$.
-Definition 2.5 (VC dimension).
-The VC dimension of a partial concept class $\mathcal{H}$, denoted $\operatorname{VCdim}(\mathcal{H})$, is the largest integer $d$ for which there exists a size-$d$ subset of $\mathcal{X}$ that is shattered by $\mathcal{H}$.
-If no such maximum $d$ exists, we define $\operatorname{VCdim}(\mathcal{H})=\infty$.
-Viewed as a matrix, the VC dimension of $\mathcal{H}$ is the largest $d$ such that the associated partial matrix $M_{\mathcal{X}\times\mathcal{H}}$ contains a $d\times 2^{d}$ submatrix with ±$\pm 1$ entries, where the columns enumerate all $d$-bit ±$\pm 1$ patterns.
-The fundamental theorem of PAC learning asserts that a total concept class is PAC learnable if and only if its VC dimension is finite.
-This result was extended to partial concept classes by [AHHM22] using a completely different proof.
-Theorem 2.6 ([AHHM22]).
-The following statements are equivalent for any partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$.
-  • $\operatorname{VCdim}(\mathcal{H})<\infty$;
-  • $\mathcal{H}$ is PAC learnable;
-Littlestone dimension:
-In his influential work, Littlestone [Lit88] provided a combinatorial characterization of concept classes that are online learnable.
-To this end, he introduced a measure that extends the VC dimension by considering the shattering of decision trees rather than sets.
-A mistake tree of depth $d$ over a domain $\mathcal{X}$ is a complete binary tree of depth $d$ with the following properties:
-  • Each internal node is labeled by an element $x\in\mathcal{X}$.
-  • Each edge is labeled by a binary value $b\in\left\{\pm 1\right\}$, where $b=-1$ indicates a left child and $b=1$ indicates a right child.
-Every root-to-leaf path in the tree is described by a sequence
-$(x_{1},b_{1}),\ldots,(x_{d},b_{d}),$
-where $x_{i}\in\mathcal{X}$ is the label of the $i$th internal node along the path, and $b_{i}$ specifies whether the path moves to the left or right child at each level.
-A concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ shatters a mistake tree if, for every root-to-leaf path
-$(x_{1},b_{1}),\ldots,(x_{d},b_{d}),$
-there exists a hypothesis $h\in\mathcal{H}$ such that $h(x_{i})=b_{i}$ for all $i\in[d]$.
-Definition 2.7 (Littlestone Dimension).
-The Littlestone dimension of a concept class $\mathcal{H}$, denoted $\operatorname{Ldim}(\mathcal{H})$, is the largest integer $d$ for which there exists a mistake tree of depth $d$ that is shattered by $\mathcal{H}$.
-If no such maximum $d$ exists, we define $\operatorname{Ldim}(\mathcal{H})=\infty$.
-It always holds that $\operatorname{VCdim}(\mathcal{H})\leq\operatorname{Ldim}(\mathcal{H})$, since any set $C=\left\{x_{1},\ldots,x_{d}\right\}$ shattered by $\mathcal{H}$ gives rise to a mistake tree of depth $d$, where all nodes at level $i$ are labeled with $x_{i}$.
-This tree is shattered by $\mathcal{H}$.
-Littlestone proved that a total concept class $\mathcal{H}$ is online learnable if and only if $\operatorname{Ldim}(\mathcal{H})<\infty$.
-This result was later extended to partial concept classes by [AHHM22].
-2.2 Differentially Private Learnability
-The widely adopted approach for ensuring privacy in machine learning is the differential privacy (DP) framework, introduced in [DMNS06].
-Informally, differential privacy in learning means that no single labeled example in the input dataset significantly impacts the learner’s output hypothesis.
-In other words, the output distribution of a differentially private randomized learning algorithm remains nearly unchanged if a single data point is modified.
-Differential privacy is quantified with two parameters $\epsilon,\delta>0$.
-We say that two probability distributions $p$ and $q$ are $(\epsilon,\delta)$-indistinguishable, if for every event $E$, we have
-$p(E)\leq e^{\epsilon}q(E)+\delta\ \text{ and }\ q(E)\leq e^{\epsilon}p(E)+\delta.$
-Two random variables are $(\epsilon,\delta)$-indistinguishable if their distributions satisfy this condition.
-Definition 2.8 (Differential Privacy).
-Given $\epsilon,\delta>0$, a randomized learning rule
-$\bm{\mathcal{A}}:(\mathcal{X}\times\left\{\pm 1\right\})^{n}\to\left\{\pm\right\}^{\mathcal{X}}$
-is $(\epsilon,\delta)$-differentially-private if for every two samples $S,S^{\prime}\in(\mathcal{X}\times\left\{\pm\right\})^{n}$ differing on a single example, the random variables $\bm{\mathcal{A}}(S)$ and $\bm{\mathcal{A}}(S^{\prime})$ are $(\epsilon,\delta)$-indistinguishable.
-We emphasize that $(\epsilon,\delta)$-indistinguishability must hold for every such pair of samples, regardless of whether they are drawn from a (realizable) distribution.
-The special case where $\delta=0$ is known as pure differential privacy, while the more general case where $\delta>0$ is referred to as approximate differential privacy.
-In approximate differential privacy, the parameters $\epsilon$ and $\delta$ are typically set as follows: $\epsilon$ is taken to be a small fixed constant (e.g., 0.1), while $\delta$ is a negligible function, $\delta=n^{\omega(1)}$, where $n$ is the sample size.
-Definitions 2.9 and 2.10 formally define differentially private learnability under the assumption that the privacy parameter $\epsilon$ is fixed at $\epsilon=0.1$.
-Definition 2.9 (Approximate Differentially Private Learnability).
-We say that a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is approximate differentially private learnable (DP-learnable) if there is a learning rule $\bm{\mathcal{A}}:(\mathcal{X}\times\left\{\pm 1\right\})^{*}\to\left\{\pm 1\right\}^{\mathcal{X}}$ with sample complexity $n(\epsilon,\delta,\delta^{\prime})$ such that the following holds.
-  • For every $\epsilon,\delta,\delta^{\prime}>0$, the class $\mathcal{H}$ is $(\epsilon,\delta)$-PAC learnable by $\bm{\mathcal{A}}$ using $n(\epsilon,\delta,\delta^{\prime})$ samples.
-  • For every $\epsilon,\delta,\delta^{\prime}>0$, the learning rule $\bm{\mathcal{A}}$ applied to samples of size $n(\epsilon,\delta,\delta^{\prime})$ is $(0.1,\delta^{\prime})$-differentially private.
-  • For every fixed $\epsilon,\delta>0$, we have
-    $\lim_{\delta^{\prime}\to 0}\frac{\log(1/\delta^{\prime})}{\log n(\epsilon,\delta,\delta^{\prime})}=\infty.$
-Definition 2.10 (Pure Differentially Private Learnability).
-We say that a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is pure differentially private learnable (pure DP-learnable) if $\mathcal{H}$ is PAC learnable by a $(0.1,0)$-differentially private learning rule.
-3 Proof of the Main Theorem
+    Question 4 in [AHHM21] also asks if every disambiguation of $\mathcal{H}_{\gamma}^{d}$ satisfies $\operatorname{VCdim}=\omega(1)$.
+Chornomaz, Moran, and Waknine explored this problem using a topological approach, but the question remains open  [CMW25].
+3 Proof of Theorem 2.1
 3.1 The lower bound
@@
 We prove that $L\geq\frac{d}{2}+1$.
-By the definition of list replicability, for any $\delta>0$, there is an integer $n$ so that for any realizable distribution $\mathcal{D}$, there exists a list of hypotheses $\{h_{1},\dots,h_{L}\}$ with
-$\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\mathbf{S})\in\{h_{1},\dots,h_{L}\}]\geq 1-\delta\text{ and }\mathcal{L}_{\mathcal{D}}(h_{i})\leq\epsilon\text{ for all }i\in[L].$
+By the definition of list replicability, for any $\delta>0$, there is an integer $n$ so that for any realizable distribution $\mu$, there exists a list of hypotheses $\{h_{1},\dots,h_{L}\}$ with
+$\operatorname{Pr}_{\mathbf{S}\sim\mu^{n}}[\bm{\mathcal{A}}(\mathbf{S})\in\{h_{1},\dots,h_{L}\}]\geq 1-\delta\text{ and }\operatorname{loss}_{\mu}(h_{i})\leq\epsilon\text{ for all }i\in[L].$
 Now pick any $\alpha>0$ and $\epsilon^{\prime}\in(\epsilon,1/2)$.
-By taking $\delta$ sufficiently small, for any distribution $\mathcal{D}$, we can choose a hypothesis $h_{\mathcal{D}}\in\{h_{1},\dots,h_{L}\}$ such that
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{\mathcal{D}}]>\frac{1}{L+\alpha}\text{ and }\mathcal{L}_{\mathcal{D}}(h_{\mathcal{D}})<\epsilon^{\prime}.$ (3)
-We will focus on a certain collection of realizable distributions $\mathcal{D}$ on $\mathbb{S}^{d-1}\times\{\pm 1\}$.
-For any $w\in\mathbb{S}^{d-1}$, take $\mathcal{D}_{w}$ to be the uniform distribution on the set $\{(x,h_{w}(x))\ |\ x\in{\mathrm{supp}}(h_{w})\}$.
+By taking $\delta$ sufficiently small, for any distribution $\mu$, we can choose a hypothesis $h_{\mu}\in\{h_{1},\dots,h_{L}\}$ such that
+$\Pr_{\mathbf{S}\sim\mu^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{\mu}]>\frac{1}{L+\alpha}\text{ and }\operatorname{loss}_{\mu}(h_{\mu})<\epsilon^{\prime}.$ (4)
+We will focus on a certain collection of realizable distributions $\mu$ on $\mathbb{S}^{d-1}\times\{\pm 1\}$.
+For any $w\in\mathbb{S}^{d-1}$, take $\mu_{w}$ to be the uniform distribution on the set $\{(x,c_{w}(x))\ |\ x\in{\mathrm{supp}}(c_{w})\}$.
 These distributions are, by definition, realizable.
-Hence, for each $\mathcal{D}_{w}$, we can choose some particular hypothesis $h_{\mathcal{D}_{w}}$ that satisfies the conditions in Eq.
-3.
+Hence, for each $\mu_{w}$, we can choose some particular hypothesis $h_{\mu_{w}}$ that satisfies the conditions in (4).
 Collect these hypotheses in a set $T$, that is to say
-$T\coloneqq\{h_{\mathcal{D}_{w}}\ |\ w\in\mathbb{S}^{d-1}\}.$
-For each $h\in T$, define the set $C_{h}\subset\mathbb{S}^{d-1}$ as
-$C_{h}\coloneqq\left\{w\in\mathbb{S}^{d-1}\ |\ \Pr_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h]>\frac{1}{L+\alpha}\text{ and }\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}\right\}.$
+$T\coloneqq\{h_{\mu_{w}}\ |\ w\in\mathbb{S}^{d-1}\}.$
+For each $h\in T$, define the set $U_{h}\subset\mathbb{S}^{d-1}$ as
+$U_{h}\coloneqq\left\{w\in\mathbb{S}^{d-1}\ |\ \Pr_{\mathbf{S}\sim\mu_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h]>\frac{1}{L+\alpha}\text{ and }\operatorname{loss}_{\mu_{w}}(h)<\epsilon^{\prime}\right\}.$
 Claim 3.2.
-The family $\{C_{h}\}_{h\in T}$ forms an antipodal-free open cover of $\mathbb{S}^{d-1}$.
+The family $\{U_{h}\}_{h\in T}$ forms an antipodal-free open cover of $\mathbb{S}^{d-1}$.
 Proof.
-The fact that any set $C_{h}$ is antipodal-free follows from the accuracy constraint $\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}$.
-Indeed, for any $w\in\mathbb{S}^{d-1}$, the concepts $h_{w}$ and $h_{-w}$ have identical support, on which they disagree at every point.
+The fact that any set $U_{h}$ is antipodal-free follows from the accuracy constraint $\operatorname{loss}_{\mu_{w}}(h)<\epsilon^{\prime}$.
+Indeed, for any $w\in\mathbb{S}^{d-1}$, the concepts $c_{w}$ and $c_{-w}$ have identical support, on which they disagree at every point.
 Thus the population loss of any hypothesis $h$ satisfies the equation
-$\mathcal{L}_{\mathcal{D}_{w}}(h)+\mathcal{L}_{\mathcal{D}_{-w}}(h)=1.$
-For any $w\in C_{h}$, we have that $\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}<1/2$, whereby $w$ and $-w$ cannot both be in $C_{h}$.
-Next, each set $C_{h}$ is open because both $\Pr_{\bm{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\bm{S})=h]$ and $\mathcal{L}_{\mathcal{D}_{w}}(h)$ are continuous in $w$.
-Lastly, the family $\{C_{h}\}_{h\in T}$ covers $\mathbb{S}^{d-1}$ because, for any $w\in\mathbb{S}^{d-1}$, the set $C_{h_{\mathcal{D}_{w}}}$ contains $w$ by construction.
+$\operatorname{loss}_{\mu_{w}}(h)+\operatorname{loss}_{\mu_{-w}}(h)=1.$
+For any $w\in U_{h}$, we have that $\operatorname{loss}_{\mu_{w}}(h)<\epsilon^{\prime}<1/2$, whereby $w$ and w-w cannot both be in $U_{h}$.
+Next, each set $U_{h}$ is open because both $\Pr_{\bm{S}\sim\mu_{w}^{n}}[\bm{\mathcal{A}}(\bm{S})=h]$ and $\operatorname{loss}_{\mu_{w}}(h)$ are continuous in $w$.
+Lastly, the family $\{U_{h}\}_{h\in T}$ covers $\mathbb{S}^{d-1}$ because, for any $w\in\mathbb{S}^{d-1}$, the set $U_{h_{\mu_{w}}}$ contains $w$ by construction.
 ∎
-Now note that the antipodal-free open cover $\{C_{h}\}_{h\in T}$ admits a finite subcover by the compactness of the unit sphere.
-Applying Theorem 3.1 to such a finite subcover guarantees that some $w\in\mathbb{S}^{d-1}$ is contained in at least $t\coloneqq\lceil\frac{d}{2}+1\rceil$ sets $C_{h_{1}},C_{h_{2}},\dots,C_{h_{t}}$.
-Unpacking definitions reveals that the distribution $\mathcal{D}_{w}$ has the property
-$\Pr_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{1}{L+\alpha}$
-for $t$ distinct hypotheses $h_{i}\in T$. Because these $h_{i}$ are distinct, the events $[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]$ are disjoint, and therefore
-$1\geq\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}\bigcup_{i=1}^{t}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]=\sum_{i=1}^{t}\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{t}{L+\alpha}.$
+Now note that the antipodal-free open cover $\{U_{h}\}_{h\in T}$ admits a finite subcover by the compactness of the unit sphere.
+Applying Theorem 3.1 to such a finite subcover guarantees that some $w\in\mathbb{S}^{d-1}$ is contained in at least $t\coloneqq\lceil\frac{d}{2}+1\rceil$ sets $U_{h_{1}},U_{h_{2}},\dots,U_{h_{t}}$.
+Unpacking definitions reveals that the distribution $\mu_{w}$ has the property
+$\Pr_{\mathbf{S}\sim\mu_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{1}{L+\alpha}$
+for $t$ distinct hypotheses $h_{i}\in T$. Because these $h_{i}$ are distinct, the events $\bm{\mathcal{A}}(\mathbf{S})=h_{i}]$ are disjoint, and therefore
+$1\geq\operatorname{Pr}_{\mathbf{S}\sim\mu_{w}^{n}}\bigcup_{i=1}^{t}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]=\sum_{i=1}^{t}\operatorname{Pr}_{\mathbf{S}\sim\mu_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{t}{L+\alpha}.$
 It follows that $L+\alpha>t=\lceil\frac{d}{2}+1\rceil$ for any $\alpha>0$, which implies the desired lower bound $L\geq\lceil\frac{d}{2}+1\rceil$.
@@
 To prove the upper bound, we design a list replicable learning algorithm $\bm{\mathcal{A}}$ that learns $\mathcal{H}_{\gamma}^{d}$ with list size $d$ independent of $\epsilon>0$.
-Given $w\in\mathbb{S}^{d-1}$, let $\overline{h}_{w}:\mathbb{S}^{d-1}\to\{\pm 1\}$ denote the total concept class corresponding to the closed half-space defined by $w$.
-$\overline{h}_{w}(x)\coloneqq\begin{cases}1&\text{if }\langle w,x\rangle\geq 0\\ -1&\text{if }\langle w,x\rangle<0\end{cases}.$
+Given $w\in\mathbb{S}^{d-1}$, let $\overline{c}_{w}:\mathbb{S}^{d-1}\to\{\pm 1\}$ denote the total concept corresponding to the closed half-space defined by $w$.
+$\overline{c}_{w}(x)\coloneqq\begin{cases}1&\text{if }{\left\langle w,x\right\rangle}\geq 0\\ -1&\text{if }{\left\langle w,x\right\rangle}<0\end{cases}.$
 Fundamentally, as in [KKL+24], we estimate a large-margin linear separator using the average of many runs of an SVM maximum margin separator.
@@
 More formally, it is the following optimization problem over the variables $\gamma\in\mathbb{R}$ and $w\in\mathbb{S}^{d-1}$:
-$\begin{array}[]{l rll}\max&\gamma&&\\ \text{s.t. }&y_{i}{\left\langle x_{i},w\right\rangle}&\geq\gamma&\text{for }i=1,\ldots,n\\ &\gamma&\geq 0&\\ &w&\in\mathbb{S}^{d-1}&\\ \end{array}$
+[…]
 One can use semi-definite programming to solve this optimization problem efficiently—to check whether it is feasible and, if so, to find the maximizing $w$.
@@
 $y{\left\langle x,w\right\rangle}\geq\gamma\ \ \text{ for all }(x,y)\in S.$
-Furthermore, for any distribution $\mathcal{D}$ over $\mathbb{S}^{d-1}\times\{\pm 1\}$, we call $w$ a $(\gamma,\epsilon)$-separator for $\mathcal{D}$ if
-$\Pr_{(x,y)\sim\mathcal{D}}[y\langle x,w\rangle<\gamma]\leq\epsilon.$
-When learning $\mathcal{H}_{\gamma}^{d}$ in the realizable setting, for any sample set $S$ drawn from a realizable distribution $\mathcal{D}$, there is some $w$ that $\gamma$-separates $S$.
+Furthermore, for any distribution $\mu$ over $\mathbb{S}^{d-1}\times\{\pm 1\}$, we call $w$ a $(\gamma,\epsilon)$-separator for $\mu$ if
+$\Pr_{(x,y)\sim\mu}[y{\left\langle x,w\right\rangle}<\gamma]\leq\epsilon.$
+When learning $\mathcal{H}_{\gamma}^{d}$ in the realizable setting, for any sample set $S$ drawn from a realizable distribution $\mu$, there is some $w$ that $\gamma$-separates $S$.
 Therefore, the hard-SVM will be feasible and return a vector $w$ that $\gamma$-separates $S$.
-The following theorem, due to [STBWA98], says that if we take a sufficiently large sample $\bm{S}$ and compute a good separator $w$ for it using hard-SVM, then with high probability, $w$ will also be a good separator for $\mathcal{D}$.
+The following theorem, due to [STBWA98], says that if we take a sufficiently large sample $\bm{S}$ and compute a good separator $w$ for it using hard-SVM, then with high probability, $w$ will also be a good separator for $\mu$.
 Theorem 3.4 (SVM generalization bound [STBWA98, Theorem 3.5]).
-For all $\epsilon,\delta>0$, there exists $n\coloneqq n(\epsilon,\delta)$ such that the following holds. Let $\mathcal{D}$ be any distribution over $\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$.
-$\Pr_{\bm{S}\sim\mathcal{D}^{n}}\left[\text{Every $w\in\mathbb{S}^{d-1}$ that $\gamma$-separates $\bm{S}$ also $\left(\frac{\gamma}{2},\epsilon\right)$-separates $\mathcal{D}$}\right]\geq 1-\delta.$
+For all $\epsilon,\delta>0$, there exists $n\coloneqq n(\epsilon,\delta)$ such that the following holds. Let $\mu$ be any distribution over $\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$.
+$\Pr_{\bm{S}\sim\mu^{n}}\left[\text{Every $w\in\mathbb{S}^{d-1}$ that $\gamma$-separates $\bm{S}$ also $\left(\frac{\gamma}{2},\epsilon\right)$-separates $\mu$}\right]\geq 1-\delta.$
 Remark 3.5.
-To prove Theorem 3.4, one can apply [STBWA98, Theorem 3.5] to show that, with probability at least $1-\delta$, both $h_{1}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}+\frac{\gamma}{2})$ and $h_{2}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}-\frac{\gamma}{2})$ have loss at most $\frac{\epsilon}{2}$, in which case $w$ is a $\left(\frac{\gamma}{2},\epsilon\right)$-separator for $\mathcal{D}$.
+To prove Theorem 3.4, one can apply [STBWA98, Theorem 3.5] to show that, with probability at least $1-\delta$, both $h_{1}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}+\frac{\gamma}{2})$ and $h_{2}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}-\frac{\gamma}{2})$ have loss at most $\frac{\epsilon}{2}$, in which case $w$ is a $\left(\frac{\gamma}{2},\epsilon\right)$-separator for $\mu$.
 Regarding optimal bounds on $n(\epsilon,\delta)$ in Theorem 3.4, we refer the reader to [GKL20, KKL+24].
-We will also use the following simple concentration result for sums of i.i.d.
+First, we prove a simple concentration result for sums of i.i.d.
 random vectors to show that the outputs of multiple runs of hard-SVM on independent samples are typically concentrated around their mean.
 Lemma 3.6.
-Let $\bm{x}_{1},\dots,\bm{x}_{k}\in\mathbb{R}^{d}$ be i.i.d random variables with mean $\mu$ and $\|\bm{x}_{i}-\mu\|_{\infty}\leq C$. Let $\bm{Z}=\frac{1}{k}\sum_{i=1}^{k}\bm{x}_{i}$. For all $t>0$,
-$\Pr[\|\bm{Z}-\mu\|_{1}\geq t]\leq 2de^{\frac{-kt^{2}}{2d^{2}C^{2}}}.$
+[…]
+[…]
 Proof.
-We apply Hoeffding’s inequality to each coordinate and take the union bound. By Hoeffding’s inequality, for every $j\in[d]$, we have
+We apply Hoeffding’s inequality  [^4] Let $c\in\mathbb{R}$ and let $\bm{x}_{1},\ldots,\bm{x}_{n}$ be independent random variables with $\bm{x}_{i}\in[-c,c]$ and $\mathbb{E}[\bm{x}_{i}]=0$.
+For any $t>0$, $\Pr\left[\left|\sum_{i=1}^{n}\bm{x}_{i}\right|\geq t\right]\leq 2e^{-\frac{t^{2}}{2c}}.$ to each coordinate and take the union bound.
+By Hoeffding’s inequality, for every $j\in[d]$, we have
 $\Pr\left[|\bm{Z}_{j}-\mu_{j}|\geq\frac{t}{d}\right]\leq 2e^{\frac{-kt^{2}}{2d^{2}C^{2}}}.$
 Therefore, by the union bound,
-$\Pr\left[\|\bm{Z}-\mu\|_{1}\geq t\right]\leq 2de^{\frac{-kt^{2}}{2d^{2}C^{2}}}.\qed$
+[…]
 We will use a rounding scheme that ensures any small neighbourhood on $\mathbb{S}^{d-1}$ is rounded to at most $d$ points.
@@
   1. 1.
-    $\|\mathrm{round}_{\alpha}(x)-x\|_{2}<\alpha$, and
+[…]
   2. 2.
-    The set $R_{x}\coloneqq\left\{\mathrm{round}_{\alpha}(y)\ |\ y\in\mathbb{S}^{d-1}\text{and }\|x-y\|_{2}\leq\beta\right\}$ has size at most $d$.
+[…]
 Proof.
 Consider any $\frac{\alpha}{2}$-net $T$ of points in general position on $\mathbb{S}^{d-1}$ and define the rounding function as
-$\mathrm{round}_{\alpha}(x)\coloneqq\arg\min_{y\in T}\|x-y\|_{2}.$
+[…]
 Property 1 follows directly from the definition of $\mathrm{round}_{\alpha}$, so it remains to prove 2.
@@
 To see this, suppose there exists a point $y\in\mathbb{R}^{d}$ that is equidistant from each $x_{i}$, meaning there exists some $r$ such that
-$r^{2}=\|x_{i}-y\|_{2}^{2}=1+\|y\|_{2}^{2}-2{\left\langle x_{i},y\right\rangle}.$
+[…]
 Consequently, $y$ is orthogonal to the linearly independent vectors $x_{1}-x_{2}$, …, $x_{1}-x_{d+1}$, and thus $y=\vec{0}$.
 Define the map $\phi:\mathbb{S}^{d-1}\to\mathbb{R}_{\geq 0}$ as
-$\phi(x)\coloneqq\tau(x)-\min_{y\in T}\|x-y\|_{2},$
+[…]
 where $\tau(x)$ denotes the distance from $x$ to a $(d+1)$-th closest point in $T$.
@@
 We need to show that for any margin $\gamma\in(0,1)$, accuracy parameter $\epsilon\in(0,1/2)$, and dimension $d\geq 1$, we have $\operatorname{LR}_{\epsilon}(\mathcal{H}_{\gamma}^{d})\leq d$.
-We will construct a list-replicable learner that always outputs a hypothesis of the form $\overline{h}_{w}$ for some $w\in\mathbb{S}^{d-1}$.
+We will construct a list-replicable learner that always outputs a hypothesis of the form $\overline{c}_{w}$ for some $w\in\mathbb{S}^{d-1}$.
 Let $k=k(d,\gamma)$ and $n_{0}=n_{0}(\epsilon,\delta,k)$ be integers yet to be determined. Consider the following learning rule $\bm{\mathcal{A}}$ that uses the rounding scheme of Lemma 3.7.
@@
 1:for $i\leftarrow 1$ to $k$ do
-2:     Sample $\mathbf{S}_{i}\sim\mathcal{D}^{n_{0}}$.
+2:  Sample $\mathbf{S}_{i}\sim\mu^{n_{0}}$.
 3:  Let $\bm{w}_{i}\leftarrow\text{hard-SVM}(\mathbf{S}_{i})$.
 4:end for
-5:  Let $\bm{w}\leftarrow\frac{1}{k}\sum_{i=1}^{k}\bm{w}_{i}$ and $\bm{z}\leftarrow\frac{\bm{w}}{\|\bm{w}\|_{2}}$.
+[…]
 6: Let $\tilde{\bm{z}}\leftarrow\mathrm{round}_{\gamma/2}(\bm{z})$.
-7:  output the hypothesis $\overline{h}_{\tilde{\bm{z}}}$.
+7:output the hypothesis $\overline{c}_{\tilde{\bm{z}}}$.
 We first show that the learning rule $\bm{\mathcal{A}}$ presented in Algorithm 1 is a PAC learner.
@@
 Let $\bm{\mathcal{A}}$ and $\bm{w}$ be as in Algorithm 1.
-For every $\epsilon,\delta\in(0,1)$ and $k\in\mathbb{N}$, there exists $n_{0}\coloneqq n_{0}(\epsilon,\delta,k)\in\mathbb{N}$ such that for every distribution $\mathcal{D}$ realizable by $\mathcal{H}_{\gamma}^{d}$, we have
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}\|_{2}<\frac{\gamma}{2}\right]\leq\frac{\delta}{4}$ (4)
+For every $\epsilon,\delta\in(0,1)$ and $k\in\mathbb{N}$, there exists $n_{0}\coloneqq n_{0}(\epsilon,\delta,k)\in\mathbb{N}$ such that for every distribution $\mu$ realizable by $\mathcal{H}_{\gamma}^{d}$, we have
+[…]
 and
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}[\mathcal{L}_{\mathcal{D}}(\bm{\mathcal{A}}(\mathbf{S}))\geq\epsilon]\leq\frac{\delta}{4}.$ (5)
+$\Pr_{\mathbf{S}\sim\mu^{kn_{0}}}[\operatorname{loss}_{\mu}(\bm{\mathcal{A}}(\mathbf{S}))\geq\epsilon]\leq\frac{\delta}{4}.$ (6)
 Proof.
 Let $\bm{w}_{1},\dots,\bm{w}_{k}$ be as in Algorithm 1.
-Since $\mathcal{D}$ is realizable by $\mathcal{H}_{\gamma}^{d}$, for every $i$, $\bm{w}_{i}$ is a $\gamma$-separator for $\mathbf{S}_{i}$.
+Since $\mu$ is realizable by $\mathcal{H}_{\gamma}^{d}$, for every $i$, $\bm{w}_{i}$ is a $\gamma$-separator for $\mathbf{S}_{i}$.
 Therefore, by Theorem 3.4, if $n_{0}(\epsilon,\delta,k)$ is sufficiently large,
-$\Pr_{\begin{subarray}{c}\mathbf{S}_{i}\sim\mathcal{D}^{n_{0}}\end{subarray}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\right]\geq 1-\frac{\delta}{4k}.$
+$\Pr_{\begin{subarray}{c}\mathbf{S}_{i}\sim\mu^{n_{0}}\end{subarray}}\left[\Pr_{(\bm{x},\bm{y})\sim\mu}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\right]\geq 1-\frac{\delta}{4k}.$
 Thus, by the union bound,
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\text{ for all $i\in[k]$}\right]\geq 1-\frac{\delta}{4},$ (6)
+$\Pr_{\mathbf{S}\sim\mu^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mu}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\text{ for all $i\in[k]$}\right]\geq 1-\frac{\delta}{4},$ (7)
 and applying the union bound again,
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\min_{i\in[k]}\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\epsilon\right]\geq 1-\frac{\delta}{4}.$ (7)
-Finally, if $(x,y)\in\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$ satisfies $y{\left\langle x,\bm{w}_{i}\right\rangle}\geq\gamma/2$ for all $i\in[k]$, then noting that $\|\bm{w}\|\leq 1$, we have
+$\Pr_{\mathbf{S}\sim\mu^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mu}\left[\min_{i\in[k]}\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\epsilon\right]\geq 1-\frac{\delta}{4}.$ (8)
+[…]
 $y{\left\langle x,\bm{z}\right\rangle}\geq y{\left\langle x,\bm{w}\right\rangle}=y{\left\langle x,\frac{1}{k}\sum_{i=1}^{k}\bm{w}_{i}\right\rangle}\geq\gamma/2.$ (9)
-Thus, from Equation 7, we have
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}\|_{2}<\frac{\gamma}{2}\right]\leq\frac{\delta}{4}$
+Thus, from (8), we have
+[…]
 and
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{z}\right\rangle}\geq\frac{\gamma}{2}\right]\geq 1-\epsilon\right]\geq 1-\frac{\delta}{4}.$ (9)
-By applying Lemma 3.7 with $\alpha\coloneqq\gamma/2$, after rounding $\bm{z}$ to $\tilde{\bm{z}}$, we have $\|\tilde{\bm{z}}-\bm{z}\|_{2}<\frac{\gamma}{2}$.
-Thus if $(x,y)\in\mathbb{S}^{d-1}\times\{\pm 1\}$ satisfy $y{\left\langle x,\bm{z}\right\rangle}\geq\gamma/2$, then
-$y{\left\langle x,\tilde{\bm{z}}\right\rangle}=y{\left\langle x,\bm{z}\right\rangle}+y{\left\langle x,\tilde{\bm{z}}-\bm{z}\right\rangle}\geq\frac{\gamma}{2}-\|\tilde{\bm{z}}-\bm{z}\|_{2}>0,$
-namely $\overline{h}_{\tilde{\bm{z}}}(x)=y$. Thus,
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\mathcal{L}_{\mathcal{D}}(\overline{h}_{\tilde{\bm{z}}})\leq\epsilon\right]\geq 1-\frac{\delta}{4},$
+$\Pr_{\mathbf{S}\sim\mu^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mu}\left[\bm{y}{\left\langle\bm{x},\bm{z}\right\rangle}\geq\frac{\gamma}{2}\right]\geq 1-\epsilon\right]\geq 1-\frac{\delta}{4}.$ (10)
+[…]
+[…]
+namely $\overline{c}_{\tilde{\bm{z}}}(x)=y$. Thus,
+$\Pr_{\mathbf{S}\sim\mu^{kn_{0}}}\left[\operatorname{loss}_{\mu}(\overline{c}_{\tilde{\bm{z}}})\leq\epsilon\right]\geq 1-\frac{\delta}{4},$
 which completes the proof of the claim. ∎
@@
 random variables in $\mathbb{S}^{d-1}$, there exists $k=k(\gamma,d)\in\mathbb{N}$ such that
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}-\operatorname{\mathbb{E}}[\bm{w}]\|_{2}\geq\frac{\gamma\beta}{2}\right]\leq\frac{\delta}{4}.$
-Since $\bm{z}=\frac{\bm{w}}{\|\bm{w}\|_{2}}$, by applying the union bound to Eq. 4 and the above inequality, we have
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{z}-\operatorname{\mathbb{E}}[\bm{z}]\|_{2}\geq\beta\right]\leq\frac{\delta}{2}.$
-Consequently, by Lemma 3.7, with probability at least $21-\delta/$ 2, the rounding scheme ($\mathrm{round}_{\frac{\gamma}{2}}$) outputs one of at most $d$ hypotheses.
+[…]
+[…]
+[…]
+Consequently, by Lemma 3.7, with probability at least $1-\delta/2$, the rounding scheme ($\mathrm{round}_{\frac{\gamma}{2}}$) outputs one of at most $d$ hypotheses.
 Applying a union bound with 3.8 completes the proof of the upper bound of Theorem 2.1.
@@
 4 Disambiguations of gap Hamming distance
-We first prove that every disambiguation of the Gap Hamming Distance matrix family has unbounded public-coin randomized communication complexity.
-Proof of Corollary 1.15.
-We combine Theorem 1.14 with the known relationship between Littlsetone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity.
+We prove Theorem 2.5 in this section.
+The Littlestone dimension of disambiguations.
+The key to obtaining (3) is to use an embedding of bounded margin half-spaces in dimension $d$ into the Boolean cube, which allows us to disambiguate $\mathcal{H}^{d}_{\gamma}$ using a disambiguation of the Gap Hamming Distance problem in dimension $O(d)$.
+The existence of such an embedding was proved in [HHM23], which we rephrase as follows.
+Lemma 4.1 (Adapted from [HHM23, Lemma 3.2]).
+Let $\gamma\in(0,1)$ and $n\in\mathbb{N}$.
+There exist $d=\Omega\left((1-\gamma)^{2}\cdot n/\log(1/(1-\gamma))\right)$, $\gamma^{\prime}\in(0,1)$ and a map $\xi\colon\mathbb{S}^{d-1}\to\left\{\pm 1\right\}^{n}$ such that for all $u,v\in\mathbb{S}^{d-1}$, we have
+${\left\langle u,v\right\rangle}\geq\gamma^{\prime}\implies{\left\langle\xi(u),\xi(v)\right\rangle}\geq\gamma n$
+${\left\langle u,v\right\rangle}\leq-\gamma^{\prime}\implies{\left\langle\xi(u),\xi(v)\right\rangle}\leq-\gamma n$
+Fix $\gamma\in(0,1)$.
+Let $\{M_{n}\}_{n=1}^{\infty}$ be a family of total functions which disambiguates $\{\operatorname{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$, and let $d=d(n)$ and $\gamma^{\prime}$ be as provided by Lemma 4.1.
+We will use this lemma along with the functions $\left\{M_{n}\right\}_{n=1}^{\infty}$ to disambiguate the family of partial concept classes $\{\mathcal{H}_{\gamma^{\prime}}^{d(n)}\}_{n=1}^{\infty}$.
+To this end, we disambiguate each partial concept $c_{w}\in\mathcal{H}_{\gamma^{\prime}}^{d}$ (defined in (1)) to
+$\overline{c}_{w}(x)\coloneqq M_{n}\big(\xi(w),\xi(x)\big).$
+Let us verify that $\overline{c}_{w}$ is, in fact, a disambiguation of $c_{w}$.
+Suppose that $c_{w}(x)=1$. By definition, this occurs exactly when ${\left\langle w,x\right\rangle}\geq\gamma^{\prime}$. It follows from the properties of $\xi$ that
+${\left\langle\xi(w),\xi(x)\right\rangle}\geq\gamma n.$
+Therefore, for such $w,x$, we have
+$\overline{c}_{w}(x)=M_{n}(\xi(w),\xi(x))=\operatorname{GHD}_{\gamma}^{n}(\xi(w),\xi(x))=1=c_{w}(x).$
+A similar argument shows that if $c_{w}(x)=-1$, then $\overline{c}_{w}(x)=-1$.
+We deduce that $\overline{c}_{w}$ indeed disambiguates $c_{w}$, and $\overline{\mathcal{H}}^{d}_{\gamma^{\prime}}$ is a disambiguation of $\mathcal{H}^{d}_{\gamma^{\prime}}$.
+Finally, note that by construction, any shattered mistake tree in $\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}$ corresponds to a shattered mistake tree of the same depth in $M_{n}$.
+Therefore, $\operatorname{Ldim}(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d})\leq\operatorname{Ldim}(M_{n})$.
+This combined with Theorem 2.3 implies that
+$\operatorname{Ldim}(M_{n})\geq\operatorname{Ldim}\left(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}\right)=\Omega\left(\sqrt{\log d(n)}\right)=\Omega\left(\sqrt{\log n}\right).$
+Communication complexity of disambiguations.
+To complete the proof of Theorem 2.5 we use the known relationships between Littlestone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity.
 Given a matrix $M\in\{\pm 1\}^{\mathcal{X}\times\mathcal{Y}}$, the margin of $M$ is defined
@@
 Let $\{M_{n}\}_{n=1}^{\infty}$ be a family of disambiguation of $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$.
-By Theorem 1.14, we know that $\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n})$.
-It thus follows from LABEL:{Ldim_of_margin} that
+By (3), we know that $\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n})$.
+It thus follows from (2) that
 $\operatorname{m}(M_{n})=O\left(\frac{1}{\sqrt[4]{\log n}}\right).$
 Finally, invoking the equivalence of margin and discrepancy by [LS09] and the relation between discrepancy and randomized communication complexity by [CG88] (see also [HHP+22, Proposition 3.3 ]) shows that the public-coin randomized communication complexity of $M_{n}$ is
-$\Omega(\log(\operatorname{m}(M_{n})^{-1}))=\Omega(\log\log n).\qed$
-We next present the proof of Theorem 1.14.
-The key is to use an embedding of bounded margin half-spaces in dimension $d$ into the Boolean cube, which allows us to disambiguate $\mathcal{H}^{d}_{\gamma}$ using a disambiguation of the Gap Hamming Distance problem in dimension $O(d)$.
-The existence of such an embedding was proved in [HHM23], which we rephrase as follows.
-Lemma 4.1 (Adapted from [HHM23, Lemma 3.2]).
-Let $\gamma\in(0,1)$ and $n\in\mathbb{N}$.
-There exist $d=\Omega\left((1-\gamma)^{2}\cdot n/\log(1/(1-\gamma))\right)$, $\gamma^{\prime}\in(0,1)$ and a map $\xi\colon\mathbb{S}^{d-1}\to\left\{\pm 1\right\}^{n}$ such that for all $u,v\in\mathbb{S}^{d-1}$, we have
-$\langle u,v\rangle\geq\gamma^{\prime}\implies\langle\xi(u),\xi(v)\rangle\geq\gamma n$
-$\langle u,v\rangle\leq-\gamma^{\prime}\implies\langle\xi(u),\xi(v)\rangle\leq-\gamma n$
-We are now ready to prove Theorem 1.14.
-Proof of Theorem 1.14.
-Fix $\gamma\in(0,1)$.
-Let $\{M_{n}\}_{n=1}^{\infty}$ be a family of total functions which disambiguates $\{\operatorname{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$, and let $d=d(n)$ and $\gamma^{\prime}$ be as provided by Lemma 4.1.
-We will use this lemma along with the functions $\left\{M_{n}\right\}_{n=1}^{\infty}$ to disambiguate the family of partial concept classes $\{\mathcal{H}_{\gamma^{\prime}}^{d(n)}\}_{n=1}^{\infty}$.
-To this end, we disambiguate each partial concept $h_{w}\in\mathcal{H}_{\gamma^{\prime}}^{d}$ (defined in LABEL:{eq:def_large_margin}) to
-$\overline{h}_{w}(x)\coloneqq M_{n}\big{(}\xi(w),\xi(x)\big{)}.$
-Let us verify that $\overline{h}_{w}$ is, in fact, a disambiguation of $h_{w}$.
-Suppose that $h_{w}(x)=1$. By definition, this occurs exactly when $\langle w,x\rangle\geq\gamma^{\prime}$. It follows from the properties of $\xi$ that
-$\langle\xi(w),\xi(x)\rangle\geq\gamma n.$
-Therefore, for such $w,x$, we have
-$\overline{h}_{w}(x)=M_{n}(\xi(w),\xi(x))=\operatorname{GHD}_{\gamma}^{n}(\xi(w),\xi(x))=1=h_{w}(x).$
-A similar argument shows that if $h_{w}(x)=-1$, then $\overline{h}_{w}(x)=-1$.
-We deduce that $\overline{h}_{w}$ indeed disambiguates $h_{w}$, and ¯$\overline{\mathcal{H}}^{d}_{\gamma^{\prime}}$ is a disambiguation of $\mathcal{H}^{d}_{\gamma^{\prime}}$.
-Finally, note that by construction, any shattered mistake tree in ¯$\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}$ corresponds to a shattered mistake tree of the same depth in $M_{n}$.
-Therefore, $\operatorname{Ldim}(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d})\leq\operatorname{Ldim}(M_{n})$.
-This combined with Theorem 1.8 implies that
-$\operatorname{Ldim}(M_{n})\geq\operatorname{Ldim}\left(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}\right)=\Omega\left(\sqrt{\log d(n)}\right)=\Omega\left(\sqrt{\log n}\right).\qed$
-Appendix A Equivalence of Stability and List-Replicability
-We prove the following more refined version of Theorem 1.2.
-Theorem A.1.
-Let $\mathcal{H}$ be any total or partial concept class on the domain $\mathcal{X}$. Then for every $\epsilon\in(0,1)$,
-$\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}\qquad\text{ and }\qquad\operatorname{LR}_{\epsilon}(\mathcal{H})\leq\frac{1}{\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{H})}.$
-Consequently, $\rho^{\mathrm{gs}}(\mathcal{H})=\frac{1}{\operatorname{LR}(\mathcal{H})}$.
+$\Omega(\log(\operatorname{m}(M_{n})^{-1}))=\Omega(\log\log n).$
+Acknowledgments
+We are grateful to Arkadev Chattopadhyay for pointing out the connection to pseudo-determinism and for valuable comments on the exposition.
+We also thank Zachary Chase for clarifying the implicit bound in [GGKM21].
+Appendix A Replicability and privacy notions
+In this section, we state the formal definitions of shared-randomness replicability, DP-learnability, and global stability.
+A.1 Shared-randomness replicability
+Let $\mathcal{A}(S,r)$ be a randomized learning rule, where $r$ denotes the random seed.
+Definition A.1 (Shared-randomness replicability [GKM21, ILPS22]).
+A concept class $\mathcal{C}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$ is shared-randomness replicable if there exists a learning rule $\mathcal{A}$ and a sample complexity function $n(\epsilon,\delta)$ such that, for every $\epsilon,\delta>0$ and every realizable distribution $\mu$, the following conditions hold:
+  • Small population loss: $\Pr_{\bm{S}\sim\mu^{n},\bm{r}}[\operatorname{loss}_{\mu}\left(\mathcal{A}(\bm{S},\bm{r})\right)>\epsilon]\leq\delta.$
+  • Replicability with shared randomness: $\Pr_{\bm{S,S^{\prime}}\sim\mu^{n},\bm{r}}[\mathcal{A}(\bm{S},\bm{r})=\mathcal{A}(\bm{S^{\prime}},\bm{r})]\geq 1-\delta.$
+One could consider shared-randomness replicability to be a weak form of replicability, as different executions of the algorithm can use the same random seed.
+A.2 Differential privacy
+The widely adopted approach for ensuring privacy in machine learning is the differential privacy (DP) framework, introduced in [DMNS06].
+Informally, differential privacy in learning means that no single labeled example in the input dataset significantly impacts the learner’s output hypothesis.
+In other words, the output distribution of a differentially private randomized learning algorithm remains nearly unchanged if a single data point is modified.
+Differential privacy is quantified with two parameters $\epsilon,\delta>0$.
+We say that two probability distributions $p$ and $q$ are $(\epsilon,\delta)$-indistinguishable, if for every event $E$, we have
+[…]
+Two random variables are $(\epsilon,\delta)$-indistinguishable if their distributions satisfy this condition.
+Definition A.2 (Differential privacy).
+Given $\epsilon,\delta>0$, a randomized learning rule
+$\bm{\mathcal{A}}:(\mathcal{X}\times\left\{\pm 1\right\})^{n}\to\left\{\pm\right\}^{\mathcal{X}}$
+is $(\epsilon,\delta)$-differentially-private if for every two samples $S,S^{\prime}\in(\mathcal{X}\times\left\{\pm\right\})^{n}$ differing on a single example, the random variables $\bm{\mathcal{A}}(S)$ and $\bm{\mathcal{A}}(S^{\prime})$ are $(\epsilon,\delta)$-indistinguishable.
+We emphasize that $(\epsilon,\delta)$-indistinguishability must hold for every such pair of samples, regardless of whether they are drawn from a (realizable) distribution.
+The special case where $\delta=0$ is known as pure differential privacy, while the more general case where $\delta>0$ is referred to as approximate differential privacy.
+In approximate differential privacy, the parameters $\epsilon$ and $\delta$ are typically set as follows: $\epsilon$ is taken to be a small fixed constant (e.g., 0.1), while $\delta$ is a negligible function, $\delta=n^{-\omega(1)}$.
+Definition A.3 (Approximate differentially drivate learnability).
+We say that a concept class $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is approximate differentially private learnable (DP-learnable) if there is a learning rule $\bm{\mathcal{A}}:(\mathcal{X}\times\left\{\pm 1\right\})^{*}\to\left\{\pm 1\right\}^{\mathcal{X}}$ with sample complexity $n(\epsilon,\delta)$ such that for every $\epsilon,\delta>0$ the following holds.
+  • The class $\mathcal{C}$ is $(\epsilon,\delta)$-PAC learnable by $\bm{\mathcal{A}}$ using $n(\epsilon,\delta)$ samples.
+  • The learning rule $\bm{\mathcal{A}}$ applied to samples of size $n(\epsilon,\delta)$ is $(\epsilon^{\prime}(n),\delta^{\prime}(n))$-differentially private learnable where $\epsilon^{\prime}(n)\leq 0.1$ and $\delta^{\prime}(n)\leq n^{-w(1)}$.
+Definition A.4 (Pure Differentially Private Learnability).
+We say that a concept class $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is pure differentially private learnable (pure DP-learnable) if $\mathcal{C}$ is PAC learnable by a $(0.1,0)$-differentially private learning rule.
+A.3 Global stability
+The concept of replicability in PAC learning first emerged in [BLM20, ABL+22] in the study of differential privacy of PAC learning algorithms.
+These works introduced a notion of replicability known as global stability to derive privacy guarantees from online learnability.
+Definition A.5.
+A learning rule $\bm{\mathcal{A}}$ for a concept class $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is $(\epsilon,\rho)$-globally stable if for every realizable distribution $\mu$, there is a hypothesis $h\in\left\{\pm 1\right\}^{\mathcal{X}}$ with population loss $\operatorname{loss}_{\mu}(h)\leq\epsilon$ satisfying
+$\Pr_{\bm{S}\sim\mu^{n}}[\bm{\mathcal{A}}(\bm{S})=h]\geq\rho,\ \ \text{ where }n=n(\epsilon).$
+We define $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{C})$ to be the supremum of $\rho$ such that there is a $(\epsilon,\rho)$-globally stable learner for $\mathcal{C}$.
+The global stability parameter of $\mathcal{C}$ is then defined as
+$\rho^{\mathrm{gs}}(\mathcal{C})\coloneqq\inf_{\epsilon>0}\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{C}).$
+The definition of global stability might initially seem weak, as a globally stable learner is not necessarily a PAC learner.
+In particular, since $\rho$ can be a small constant, there may be a probability as great as $1-\rho$ that the learning rule outputs a hypothesis with large population loss.
+However, as discussed in the next section, global stability is equivalent to the seemingly stronger notion of list replicability.
+Appendix B Equivalence of global stability and list replicability
+In [CMY23], Chase, Moran and Yehudayoff proved that for every total class $\mathcal{C}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$, list replicability is equivalent to global stability.
+It is easy to check that their proof applies to partial concept classes, resulting in the following relationship between the list replicability number and the global stability parameter.
+Theorem B.1.
+Let $\mathcal{C}$ be any total or partial concept class on the domain $\mathcal{X}$. Then for every $\epsilon\in(0,1)$,
+$\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{C})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{C})}\qquad\text{ and }\qquad\operatorname{LR}_{\epsilon}(\mathcal{C})\leq\frac{1}{\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{C})}.$
+Consequently, $\rho^{\mathrm{gs}}(\mathcal{C})=\frac{1}{\operatorname{LR}(\mathcal{C})}$.
 Proof.
-We first prove that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}$.
-Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon,L)$-list replicable learner for $\mathcal{H}$ with sample complexity $n=n(\epsilon,\delta)$.
-Let $\mathcal{D}$ be any realizable distribution on $\mathcal{X}\times\{\pm 1\}$, and let $h_{1},\ldots,h_{L}$ be the list of hypotheses guaranteed by Definition 1.1.
+We first prove that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{C})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{C})}$.
+Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon,L)$-list replicable learner for $\mathcal{C}$ with sample complexity $n=n(\epsilon,\delta)$.
+Let $\mu$ be any realizable distribution on $\mathcal{X}\times\{\pm 1\}$, and let $h_{1},\ldots,h_{L}$ be the list of hypotheses guaranteed by Definition 1.2.
 By the pigeonhole principle, at least one $h_{i}$ satisfies
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})=h_{i}]\geq\frac{1-\delta}{L}.$
+$\Pr_{\mathbf{S}\sim\mu^{n}}[\bm{\mathcal{A}}(\bm{S})=h_{i}]\geq\frac{1-\delta}{L}.$
 Since this statement holds for arbitrary $\delta>0$, $\bm{\mathcal{A}}$ is itself an $(\epsilon,\rho)$-globally stable learner for all $\rho<\frac{1}{L}$.
-We may conclude that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}$.
-Next, we prove that $\operatorname{LR}_{\epsilon}(\mathcal{H})\leq 1/\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{H})$.
-Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon/3,\rho)$-globally stable learner for $\mathcal{H}$ with sample complexity $n_{0}=n_{0}(\epsilon)$.
-By the stability assumption, for every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\{\pm 1\}$, there exists $h^{*}:\mathcal{X}\to\left\{\pm 1\right\}$ satisfying
-$\mathcal{L}_{\mathcal{D}}(h^{*})\leq\frac{\epsilon}{3}\ \text{ and }\ \Pr_{\begin{subarray}{c}\bm{S}\sim\mathcal{D}^{n_{0}}\end{subarray}}[\bm{\mathcal{A}}(\bm{S})=h^{*}]\geq\rho.$ (10)
-For every $h\in\{\pm 1\}^{\mathcal{X}}$ and realizable distribution $\mathcal{D}$, define
-$p(h)\coloneqq\Pr_{\mathbf{S}\sim\mathcal{D}^{n_{0}}}[\bm{\mathcal{A}}(\bm{S})=h],$
+We may conclude that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{C})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{C})}$.
+Next, we prove that $\operatorname{LR}_{\epsilon}(\mathcal{C})\leq 1/\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{C})$.
+Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon/3,\rho)$-globally stable learner for $\mathcal{C}$ with sample complexity $n_{0}=n_{0}(\epsilon)$.
+By the stability assumption, for every realizable distribution $\mu$ on $\mathcal{X}\times\{\pm 1\}$, there exists $h^{*}:\mathcal{X}\to\left\{\pm 1\right\}$ satisfying
+$\operatorname{loss}_{\mu}(h^{*})\leq\frac{\epsilon}{3}\ \text{ and }\ \Pr_{\begin{subarray}{c}\bm{S}\sim\mu^{n_{0}}\end{subarray}}[\bm{\mathcal{A}}(\bm{S})=h^{*}]\geq\rho.$ (11)
+For every $h\in\{\pm 1\}^{\mathcal{X}}$ and realizable distribution $\mu$, define
+$p(h)\coloneqq\Pr_{\mathbf{S}\sim\mu^{n_{0}}}[\bm{\mathcal{A}}(\bm{S})=h],$
 Denote $L\coloneqq\left\lfloor\frac{1}{\rho}\right\rfloor$, so that $\rho\in\left(\frac{1}{L+1},\frac{1}{L}\right]$, and let $\alpha\coloneqq\rho-\frac{1}{L+1}>0$.
 Define the list $\Lambda$ of good and likely hypotheses
-$\Lambda\coloneqq\left\{h\in\left\{\pm 1\right\}^{\mathcal{X}}\ |\ p(h)>\frac{1}{L+1}\text{ and }\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right\}.$
+$\Lambda\coloneqq\left\{h\in\left\{\pm 1\right\}^{\mathcal{X}}\ |\ p(h)>\frac{1}{L+1}\text{ and }\operatorname{loss}_{\mu}(h)\leq\epsilon\right\}.$
 Note that $|\Lambda|\leq L$ and $\Lambda$ is nonempty, as it contains $h^{*}$.
@@
 1: Sample a dataset:
-$\bm{S}=({\bm{P}},\bm{Q})\sim\mathcal{D}^{tn_{0}+n_{1}},\quad\text{where }{\bm{P}}=({\bm{P}}_{1},\ldots,{\bm{P}}_{t})\sim(\mathcal{D}^{n_{0}})^{t}=\mathcal{D}^{tn_{0}},\quad\text{and}\quad\bm{Q}\sim\mathcal{D}^{n_{1}}.$
+$\bm{S}=({\bm{P}},\bm{Q})\sim\mu^{tn_{0}+n_{1}},\quad\text{where }{\bm{P}}=({\bm{P}}_{1},\ldots,{\bm{P}}_{t})\sim(\mu^{n_{0}})^{t}=\mu^{tn_{0}},\quad\text{and}\quad\bm{Q}\sim\mu^{n_{1}}.$
 2: Define the empirical estimate of $p(h)$ as
@@
   • $\operatorname{freq}_{\bm{P}}(h)\geq\rho-\tfrac{\alpha}{2}$
-  • $\mathcal{L}_{{\bm{Q}}}(h)\leq\tfrac{2\epsilon}{3}$
+  • $\operatorname{loss}_{{\bm{Q}}}(h)\leq\tfrac{2\epsilon}{3}$
 If no such $h$ exists, output an arbitrary $h$ corresponding to “failure.”
@@
 To show that $\bm{\mathcal{A}}^{\prime}$ outputs a hypothesis from $\Lambda$ with probability at least $1-\delta$, we will condition on the events
-$A:\quad$ $|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}$
+$A:\quad$ $|\operatorname{loss}_{\mu}(h)-\operatorname{loss}_{\bm{Q}}(h)|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}$
 $B:\quad$ $|p(h)-\operatorname{freq}_{\bm{P}}(h)|<\frac{\alpha}{2}\text{ for all }h\in\left\{\pm 1\right\}^{\mathcal{X}}$
 To guarantee that both events are likely, we prove the following claim.
-Claim A.2.
+Claim B.2.
 There exist integers $t(\alpha,\delta)$ and $n_{1}(\epsilon,t)$ such that
-$\Pr_{{\bm{P}}\sim\mathcal{D}^{tn_{0}}}[B]\geq 1-\frac{\delta}{2}\quad\text{and}\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}[A]\geq 1-\frac{\delta}{2}.$
-Proof of A.2.
-For the choice of $t$ and the proof of the first inequality, we use the uniform convergence property of the family of indicator functions on {±$\{\pm 1\}^{\mathcal{X}}$.
+$\Pr_{{\bm{P}}\sim\mathcal{D}^{tn_{0}}}[B]\geq 1-\frac{\delta}{2}\quad\text{and}\Pr_{{\bm{Q}}\sim\mu^{n_{1}}}[A]\geq 1-\frac{\delta}{2}.$
+Proof of B.2.
+For the choice of $t$ and the proof of the first inequality, we use the uniform convergence property of the family of indicator functions on $\{\pm 1\}^{\mathcal{X}}$.
 More precisely, for $f\in\left\{\pm 1\right\}^{\mathcal{X}}$, define $\mathbb{I}_{f}:\left\{\pm 1\right\}^{\mathcal{X}}\to\left\{0,1\right\}$ as
@@
 has VC dimension $1$, and therefore, it satisfies the uniform convergence property.
-For ${\bm{P}}_{i}\sim\mathcal{D}^{n_{0}}$, $\bm{\mathcal{A}}({\bm{P}}_{i})$ induces a probability distribution $\mu$ on {±$\left\{\pm 1\right\}^{\mathcal{X}}$, and we have
-$1-p(h)=\Pr_{{\bm{P}}_{i}\sim\mathcal{D}^{n_{0}}}[\bm{\mathcal{A}}({\bm{P}}_{i})\neq h]=\mathcal{L}_{\mu}(\mathbb{I}_{h}),$
+For ${\bm{P}}_{i}\sim\mu^{n_{0}}$, $\bm{\mathcal{A}}({\bm{P}}_{i})$ induces a probability distribution $\mu$ on $\left\{\pm 1\right\}^{\mathcal{X}}$, and we have
+$1-p(h)=\Pr_{{\bm{P}}_{i}\sim\mu^{n_{0}}}[\bm{\mathcal{A}}({\bm{P}}_{i})\neq h]=\operatorname{loss}_{\mu}(\mathbb{I}_{h}),$
 while $1-\operatorname{freq}_{\bm{P}}(h)$ corresponds to the empirical loss of $\mathbb{I}_{h}$ on $(\mathbb{I}_{h_{1}},\ldots,\mathbb{I}_{h_{t}})\sim\mu^{t}$.
@@
 Now that we have $t$, we can define $n_{1}$ and prove the second inequality.
-Note that for every $h\in\{\pm 1\}^{\mathcal{X}}$, for ${\bm{Q}}\sim\mathcal{D}^{n_{1}}$, $\mathcal{L}_{\bm{Q}}(h)$ is an average of ${n_{1}}$ samplings of a Bernoulli random variable with expectation $\mathcal{L}_{\mathcal{D}}(h)$.
+Note that for every $h\in\{\pm 1\}^{\mathcal{X}}$, for ${\bm{Q}}\sim\mu^{n_{1}}$, $\operatorname{loss}_{\bm{Q}}(h)$ is an average of $n_{1}$} samplings of a Bernoulli random variable with expectation $\operatorname{loss}_{\mu}(h)$.
 Thus, by Hoeffding’s inequality, there exists $n_{1}=n_{1}(\epsilon^{\prime},t)$ such that
-$\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}\left[\left|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)\right|>\frac{\epsilon}{3}\right]\leq\frac{\delta}{2t}.$ (11)
+$\Pr_{{\bm{Q}}\sim\mu^{n_{1}}}\left[\left|\operatorname{loss}_{\mu}(h)-\operatorname{loss}_{\bm{Q}}(h)\right|>\frac{\epsilon}{3}\right]\leq\frac{\delta}{2t}.$ (12)
 Thus, by the union bound, we have
-$\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}\left[\left|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)\right|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}\right]\geq 1-\frac{\delta}{2}.$ (12)
-A direct consequence of Claim A.2 is that
-$\Pr_{\mathbf{S}\sim\mathcal{D}^{tn_{0}+n_{1}}}\left[A,B\right]\geq 1-\delta.$
-Condition on events $A$ and $B$, and let $h^{*}$ be a stable hypothesis for $\bm{\mathcal{A}}$, as described in Eq.
-10.
+$\Pr_{{\bm{Q}}\sim\mu^{n_{1}}}\left[\left|\operatorname{loss}_{\mu}(h)-\operatorname{loss}_{\bm{Q}}(h)\right|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}\right]\geq 1-\frac{\delta}{2}.$ (13)
+A direct consequence of Claim B.2 is that
+$\Pr_{\mathbf{S}\sim\mu^{tn_{0}+n_{1}}}\left[A,B\right]\geq 1-\delta.$
+Condition on events $A$ and $B$, and let $h^{*}$ be a stable hypothesis for $\bm{\mathcal{A}}$, as described in (11).
 We will show that $h^{*}$ is a candidate for output, so $\bm{\mathcal{A}}^{\prime}$ will not output “failure”.
-To check the first condition for output, we combine $B$ and Eq.
-10 to show that
+To check the first condition for output, we combine $B$ and (11) to show that
 $\operatorname{freq}_{\bm{P}}(h^{*})\geq p(h^{*})-\frac{\alpha}{2}\geq\rho-\frac{\alpha}{2}.$
 Moreover, $\rho-\frac{\alpha}{2}>0$, so $h^{*}\in\mathcal{Y}$. We may therefore apply $A$ to show that $h^{*}$ satisfies the second condition for output,
-$\mathcal{L}_{\bm{Q}}(h^{*})\leq\mathcal{L}_{\mathcal{D}}(h^{*})+\frac{\epsilon}{3}\leq\frac{2\epsilon}{3}.$
+$\operatorname{loss}_{\bm{Q}}(h^{*})\leq\operatorname{loss}_{\mu}(h^{*})+\frac{\epsilon}{3}\leq\frac{2\epsilon}{3}.$
 Finally, let $h_{o}$ be any output of $\bm{\mathcal{A}}^{\prime}$, conditioned on $A$ and $B$.
@@
 $p(h_{o})>\operatorname{freq}_{\bm{P}}(h_{o})-\frac{\alpha}{2}\geq\rho-\alpha=\frac{1}{L+1}.$
-Furthermore, $h_{o}$ also satisfies the condition $\mathcal{L}_{\bm{Q}}(h_{o})\leq\frac{2\epsilon}{3}$, so because of $A$,
-$\mathcal{L}_{\mathcal{D}}(h_{o})\leq\mathcal{L}_{\bm{Q}}(h_{o})+\frac{\epsilon}{3}\leq\epsilon.$
+Furthermore, $h_{o}$ also satisfies the condition $\operatorname{loss}_{\bm{Q}}(h_{o})\leq\frac{2\epsilon}{3}$, so because of $A$,
+$\operatorname{loss}_{\mu}(h_{o})\leq\operatorname{loss}_{\bm{Q}}(h_{o})+\frac{\epsilon}{3}\leq\epsilon.$
 Thus, $h_{o}$ must be in $\Lambda$. ∎